A Complete Bibliography of ECMA Standards

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 April 2024 Version 1.09

Title word cross-reference

128 [ECM97-37]. 3 [ECM04q]. 7 [ECM91e]. 8 [ECM99a, ECM00a, ECM86a, ECM00b, ECM99b, ECM00c, ECM91c, ECM86b].

-Bit [ECM99a, ECM00a, ECM86a, ECM00b, ECM99b, ECM00c, ECM91c, ECM91e, ECM86b]. -Track [ECM97-37].

/CLI [ECM05j].

1 [ECM95d, ECM96i, ECM98f, ECM98g, ECM98h, ECM99n, ECM00-31, ECM01a, ECM04j, ECM04p]. 100 [ECM88]. 106 [ECM93a]. 107 [ECM95a]. 108 [ECM96a]. 109 [ECM96b]. 112-Track [ECM93j]. 113 [ECM99a]. 114 [ECM00a]. 118 [ECM86a]. 119 [ECM87a]. 120 [ECM93b, ECM02a].

121 [ECM00b]. 125 [ECM87b]. 128 [ECM99b]. 128-Track [ECM94g, ECM95h, ECM00g]. 13 [ECM85a]. 130 [ECM96c]. 133 [ECM98a]. 139 [ECM90a]. 142 [ECM97a]. 143 [ECM97b]. 144 [ECM00c]. 145 [ECM90b]. 146 [ECM90c]. 147 [ECM90d]. 148 [ECM97c]. 149 [ECM97d]. 150 [ECM92a]. 151 [ECM91a]. 152 [ECM93c]. 153 [ECM94a]. 154 [ECM94b]. 155 [ECM97e]. 156 [ECM93d]. 157 [ECM93e]. 158 [ECM97f]. 159 [ECM91b]. 160 [ECM92b]. 161 [ECM93f]. 162 [ECM97g]. 163 [ECM97h]. 164 [ECM97i]. 165 [ECM97j]. 167 [ECM97k]. 168 [ECM94c]. 169 [ECM92c]. 170 [ECM92d]. 171 [ECM92e]. 172 [ECM92f]. 173 [ECM97l]. 174 [ECM97m]. 175 [ECM98b]. 176 [ECM98c]. 177 [ECM97n]. 178 [ECM97o]. 179 [ECM92g]. 18-Track [ECM93b, ECM93c].

180 [ECM92h]. 182 [ECM92i]. 183 [ECM92j]. 184 [ECM92k]. 185 [ECM97p]. 186 [ECM00d]. 189 [ECM93g]. 190 [ECM93h]. 191 [ECM97q]. 192 [ECM97r]. 193 [ECM97s]. 194 [ECM97t]. 195 [ECM95b]. 196 [ECM93i]. 197 [ECM93j]. 198 [ECM95c]. 199 [ECM93k].

2 [ECM95c, ECM98i, ECM99o, ECM99p, ECM00-30]. 200 [ECM93l]. 201 [ECM94d]. 202 [ECM97u]. 203 [ECM97v]. 205 [ECM93m]. 206 [ECM93n]. 207 [ECM94e]. 208 [ECM94f]. 208-Track [ECM97-38, ECM00m]. 209 [ECM94g]. 21 [ECM94e]. 210 [ECM95d]. 211 [ECM97w]. 212 [ECM97x]. 213 [ECM97y]. 214 [ECM97z]. 215 [ECM97-27]. 216 [ECM97-28]. 217 [ECM94h]. 218 [ECM94i]. 219 [ECM96d]. 220 [ECM97-29]. 221 [ECM97-30]. 222 [ECM95e]. 223 [ECM95f]. 224 [ECM97-31]. 225 [ECM97-32]. 226 [ECM95g]. 230 [ECM94d, ECM97-33]. 231 [ECM95h]. 232 [ECM95i]. 233 [ECM97-34]. 234 [ECM95j]. 235 [ECM96e]. 236 [ECM96f]. 237 [ECM96g]. 238 [ECM96h]. 239 [ECM96i]. 240 [ECM96j]. 241 [ECM98d]. 242 [ECM98e]. 243 [ECM97-35]. 244 [ECM96k]. 245 [ECM97-36]. 246 [ECM98f]. 247 [ECM98g]. 248 [ECM98h]. 249 [ECM98i]. 250 [ECM98j]. 251 [ECM98k]. 252 [ECM96l]. 253 [ECM00e]. 254 [ECM99c]. 258 [ECM97-37]. 259 [ECM97-38]. 260 [ECM97-39]. 261 [ECM97-40]. 262 [ECM99d]. 263 [ECM98l]. 264 [ECM98m]. 265 [ECM97-41]. 266 [ECM97-42]. 267 [ECM99e]. 268 [ECM99f]. 269 [ECM00f]. 270 [ECM97-43]. 271 [ECM99g]. 272 [ECM99h]. 273 [ECM98n]. 274 [ECM99i]. 275 [ECM98o]. 276 [ECM98p]. 277 [ECM98q]. 278 [ECM00g]. 279 [ECM98r]. 280 [ECM98s]. 281 [ECM00h]. 282 [ECM00i]. 283 [ECM00j]. 284 [ECM00k]. 285 [ECM00l]. 286 [ECM00m]. 287 [ECM99j]. 288 [ECM99k].

289 [ECM99l]. 290 [ECM99m]. 291

[ECM99n]. 292 [ECM99o]. 293 [ECM99p]. 294 [ECM99q]. 295 [ECM99r]. 296 [ECM99s]. 297 [ECM99t]. 298 [ECM99u]. 299 [ECM00n].

3 [ECM96f, ECM97-37, ECM01h]. 3-XT [ECM97-37]. 300 [ECM00o]. 301 [ECM00p]. 302 [ECM00q]. 303 [ECM00r]. 304 [ECM00s]. 305 [ECM00t]. 306 [ECM00u]. 307 [ECM00v]. 308 [ECM00w]. 309 [ECM00x]. 310 [ECM00y]. 311 [ECM00z]. 312 [ECM00-27]. 313 [ECM00-28]. 314 [ECM00-29]. 315 [ECM00-30]. 316 [ECM00-31]. 317 [ECM00-32]. 318 [ECM00-33]. 319 [ECM01a]. 320 [ECM01b]. 321 [ECM01c]. 322 [ECM01d]. 323 [ECM04a]. 324 [ECM01e]. 325 [ECM01f]. 326 [ECM01g]. 327 [ECM01h]. 328 [ECM01i]. 329 [ECM01j]. 330 [ECM02a]. 331 [ECM04b]. 332 [ECM01k]. 333 [ECM03a]. 334 [ECM01l, ECM02b, ECM05a, ECM06a]. 335 [ECM01m, ECM02c, ECM05b, ECM06b]. 336 [ECM02d]. 337 [ECM03b]. 338 [ECM03c]. 339 [ECM02e]. 340 [ECM04c]. 341 [ECM04d]. 342 [ECM02f]. 343 [ECM03d]. 344 [ECM03e]. 345 [ECM03f]. 346 [ECM03g]. 347 [ECM03h]. 348 [ECM04e]. 349 [ECM04f]. 35 [ECM94j]. 350 [ECM03i]. 351 [ECM03j]. 352 [ECM03k]. 353 [ECM04g]. 354 [ECM04h]. 355 [ECM04i]. 356 [ECM04j]. 357 [ECM04k]. 358 [ECM04l]. 359 [ECM04m]. 36-Track [ECM93i]. 360 [ECM04n]. 361 [ECM04o]. 362 [ECM04p]. 363 [ECM04q]. 364 [ECM05c]. 365 [ECM05d]. 366 [ECM05e]. 367 [ECM05f]. 368 [ECM05g, ECM05h]. 369 [ECM05h]. 370 [ECM06c]. 371 [ECM05i]. 372 [ECM05j]. 373 [ECM06d]. 374 [ECM06e]. 375 [ECM06f]. 376 [ECM06g]. 377 [ECM06h]. 384-Track [ECM01a]. 385 [ECM95f].

4 [ECM99k]. 400 [ECM92f]. 43 [ECM91c]. 448-Track [ECM01b]. 48 [ECM91d]. 48-Track [ECM92i]. 4X [ECM03b, ECM05c, ECM06e].

5 [ECM92f].

6 [ECM91e]. 650 [ECM96i, ECM96j].

74 [ECM99v].

8 [ECM01j]. 80 [ECM03c, ECM04m]. 8X [ECM04f, ECM05i].

94 [ECM86b]. 99 [ECM85b].

Abstract [ECM97d]. Access [ECM98q, ECM03f]. Ada [ECM97g]. Adaptation [ECM97-41]. Adaptive [ECM91a, ECM95e]. Additional [ECM98b, ECM98c, ECM97-27, ECM97-29, ECM97-30, ECM97-31, ECM97-32, ECM97-34, ECM98j, ECM98k, ECM96l, ECM00j, ECM00k, ECM00p, ECM00q, ECM00r, ECM00s, ECM00-28, ECM00-29]. Addressing [ECM97e]. Advice [ECM97x, ECM97w]. Air [ECM00-27, ECM98o]. Airborne [ECM99v]. AIT [ECM01j, ECM98f, ECM99n, ECM99o]. AIT-1 [ECM98f, ECM99n]. AIT-2 [ECM99o]. AIT-3 [ECM01j]. Algorithm [ECM91a, ECM91b, ECM95e, ECM01c]. Alphabet [ECM99a, ECM00a, ECM86a, ECM00b, ECM99b, ECM00c]. Alphabets [ECM86b]. AMDs [ECM98o]. analysis [ECM05f]. ANF [ECM98b, ECM97-29, ECM97-31, ECM98j]. ANF-CINTSD [ECM97-29]. ANF-CMNSD [ECM98j]. ANF-PRSD [ECM98b]. ANF-TCSD [ECM97-31]. AOC [ECM97x]. AOCSD [ECM97w]. API [ECM96e]. APIW [ECM95j]. Application [ECM96d, ECM95j, ECM04h, ECM05e]. Applications [ECM04a, ECM92g, ECM92h, ECM00f, ECM00l, ECM94h, ECM94i]. Arabic [ECM00a]. Arithmetic [ECM91b]. Association [ECM93n]. ATM [ECM97-41, ECM98q]. Attribute [ECM96d]. Authentication [ECM96d, ECM97-35, ECM00t, ECM00u]. Azimuth [ECM92c].

B [ECM96l, ECM99c, ECM97-40, ECM97-41, ECM97-42, ECM99q, ECM99r, ECM99s, ECM99t, ECM99u]. B-BCSD [ECM97-40]. B-ISDN [ECM99q, ECM99r, ECM99s, ECM99t]. B-PISN [ECM96l, ECM99c, ECM97-40, ECM97-41, ECM97-42, ECM99q, ECM99r, ECM99s, ECM99t, ECM99u]. B-QSIG [ECM99q, ECM99r, ECM99s, ECM99t]. B-QSIG-BC [ECM97-42]. B-QSIG-GF [ECM99c]. B-QSIG-SAAL [ECM97-41]. B-QSIG-SBC [ECM99u]. B-QSIG-TC [ECM96l]. Basic [ECM93a, ECM96k, ECM97-42, ECM98q, ECM01k, ECM02e]. BC [ECM93a, ECM97b, ECM97-42]. BCSD [ECM97a, ECM97-40]. Bearer [ECM97a, ECM97b, ECM97-40, ECM99q, ECM99r, ECM99u]. between [ECM00v, ECM00w, ECM00x, ECM00-33, ECM01g, ECM01k, ECM02e, ECM04n, ECM04o]. Binary [ECM91b]. Binding [ECM97f, ECM97g, ECM97-33]. Bit [ECM99a, ECM00a, ECM86a, ECM00b, ECM99b, ECM00c, ECM91c, ECM91e, ECM86b]. borne [ECM98o]. Both [ECM85b]. Broadband [ECM96l, ECM99c, ECM97-40, ECM97-41, ECM97-42, ECM99q, ECM99r, ECM99s, ECM99t, ECM99u]. Business [ECM92b]. Byte [ECM99a, ECM00a, ECM86a, ECM00b, ECM99b, ECM00c, ECM86b].

C [ECM97f, ECM00-33, ECM05j]. C# [ECM01l, ECM02b, ECM05a, ECM06a]. Call [ECM97m, ECM00d, ECM97-30, ECM98m, ECM97-42, ECM99q, ECM99r, ECM00n, ECM00o, ECM00r, ECM00s, ECM00w, ECM00x, ECM00-28, ECM00-29, ECM01g, ECM03d, ECM03e, ECM04n, ECM04o, ECM97l, ECM97n, ECM97o, ECM97p, ECM97q, ECM97r, ECM97u, ECM97v, ECM97-27, ECM97-29, ECM97-34, ECM98l, ECM98m, ECM00j, ECM00k, ECM00-28]. Call/Connection [ECM97-42]. Capabilities [ECM97a]. Capacity [ECM92j, ECM92k, ECM95b, ECM94d, ECM94e, ECM95f, ECM96h, ECM96i, ECM96j, ECM97-39, ECM99i, ECM98s, ECM00-32, ECM01d, ECM03b, ECM04f, ECM03i, ECM03j, ECM04g, ECM05c, ECM05d, ECM05i, ECM06e]. Cartridge [ECM90a, ECM90b, ECM90c, ECM92a, ECM92c, ECM92d, ECM92e, ECM92j, ECM92k, ECM95b, ECM95c, ECM94d, ECM95d, ECM95f, ECM96f, ECM96h, ECM96i, ECM96j, ECM98f, ECM98g, ECM98i, ECM97-39, ECM00g, ECM98s, ECM99k, ECM99n, ECM99o, ECM99p, ECM00-31, ECM00-32, ECM01d, ECM01j, ECM03i, ECM03j, ECM04g]. Cartridges [ECM88, ECM95a, ECM93b, ECM87b, ECM90d, ECM93c, ECM94a, ECM94b, ECM92i, ECM92j, ECM92k, ECM93g, ECM93h, ECM95b, ECM93i, ECM93j, ECM94d, ECM94e, ECM94g, ECM95f, ECM95h, ECM96i, ECM96j, ECM97-37, ECM97-38, ECM97-39, ECM98s, ECM00m, ECM00-32, ECM01a, ECM01b, ECM01d, ECM03i, ECM03j, ECM04g, ECM85b]. Case [ECM98n, ECM04b, ECM06f, ECM97-43]. Cassette [ECM98h, ECM00-30]. CC [ECM00d]. CCS [ECM93h]. CCSD [ECM97p]. CD [ECM96c]. CD-ROM [ECM96c]. CDIF [ECM97-43]. CDROM [ECM87a]. Centre [ECM03f, ECM03g, ECM03h]. CF [ECM97m]. CFSD [ECM97l]. Change [ECM96j, ECM97-39]. Character [ECM99a, ECM00a, ECM86a, ECM00b, ECM99b, ECM00c, ECM94j, ECM91c, ECM91d, ECM91e, ECM86b]. Characteristics [ECM96g]. Charge [ECM97w, ECM97x]. Chemical [ECM01i]. CI [ECM97v]. CIDL [ECM00-29]. CIDLSD [ECM00-28]. CINT [ECM97-30]. CINTSD [ECM97-29]. Circuit [ECM93a, ECM97a, ECM97b, ECM95g, ECM96k, ECM00e, ECM98q, ECM99l]. CISD [ECM97u]. Class [ECM93m, ECM99g]. CLI [ECM01m, ECM02c, ECM05b, ECM06b, ECM05j]. Clock [ECM97-36]. CM [ECM95g]. CMN [ECM98k]. CMNSD [ECM98j]. CO [ECM97r]. Code [ECM94j]. Coded [ECM99a, ECM00a, ECM86a, ECM00b, ECM99b, ECM00c, ECM91c, ECM91d, ECM91e, ECM86b]. Coding [ECM91a, ECM91b]. COFC [ECM93m, ECM99g]. Commercially [ECM99g, ECM93m]. Common [ECM97d, ECM97f, ECM97g, ECM97-33, ECM98k, ECM97-43, ECM98j, ECM01m, ECM02c, ECM05b, ECM06b]. Communication [ECM04c, ECM03k, ECM06d]. Compact [ECM94c, ECM01h]. Completion [ECM01g, ECM97p, ECM00d]. Components [ECM99m]. Compression [ECM91a, ECM91b, ECM95e, ECM01c]. Computer [ECM92b, ECM92g, ECM92h, ECM00f, ECM00l, ECM04a, ECM94h, ECM94i]. Configuration [ECM98p]. Configurations [ECM98a]. Conformance [ECM99r, ECM99t]. Connection [ECM95i, ECM00e, ECM97-40, ECM97-42]. Connections [ECM95g, ECM96k, ECM99l]. considerations [ECM04d]. Context [ECM93n]. Control [ECM93f, ECM97-42, ECM99q, ECM99r, ECM99u, ECM91d]. Controlled [ECM92b]. Cordless [ECM97-27, ECM97-28, ECM97-34, ECM97-35, ECM97-27, ECM97-28, ECM97-34, ECM97-35]. Corporate [ECM00v, ECM00w, ECM00x, ECM01g, ECM01k, ECM02e, ECM04i, ECM04n, ECM04o]. COSD [ECM97q]. Counter [ECM97-31, ECM97-32, ECM96l]. CPI [ECM98l, ECM98m]. CSTA [ECM04a, ECM04e, ECM92g, ECM92h, ECM94h, ECM94i, ECM00f, ECM00l]. CT [ECM97o]. CTAU [ECM97-35]. CTLR [ECM97-28]. CTM [ECM97-27, ECM97-28, ECM97-34, ECM97-35]. CTMI [ECM97-27]. CTMO [ECM97-34]. CTSD [ECM97n]. Cyrillic [ECM99a].

D [ECM04q]. D3 [ECM95d]. DA-2 [ECM98i]. DAT-DC [ECM92e]. Data [ECM88, ECM93b, ECM87b, ECM96c, ECM90d, ECM91a, ECM93c, ECM94b, ECM91b, ECM92i, ECM92j, ECM92k, ECM95b, ECM93i, ECM93j, ECM94d, ECM94e, ECM94f, ECM94g, ECM95e, ECM95f, ECM95h, ECM96h, ECM96i, ECM96j, ECM97-37, ECM97-38, ECM97-39, ECM97-43, ECM99i, ECM00g, ECM98s, ECM00m, ECM00-32, ECM01a, ECM01b, ECM01c, ECM01d, ECM05c, ECM05d, ECM05i, ECM06e, ECM85b, ECM90c, ECM92e, ECM95d, ECM03b, ECM04f, ECM03i, ECM03j, ECM04g]. DATA-D3-1 [ECM95d]. DATA/DAT [ECM90c, ECM92e]. DATA/DAT-DC [ECM92e]. DC [ECM92a, ECM92e]. DCLZ [ECM91a]. DDS [ECM90a, ECM92a, ECM92d, ECM95c, ECM96f, ECM99k]. DDS-2 [ECM95c]. DDS-3 [ECM96f]. DDS-4 [ECM99k]. DDS-DC [ECM92a]. Declaration [ECM06c]. Declared [ECM96b]. Dedicated [ECM95g]. Definition [ECM97-33]. demand [ECM96k]. Density [ECM03i]. Description [ECM97a, ECM97-40, ECM04e]. design [ECM04d, ECM05f]. Detection [ECM01i]. Determination [ECM92b]. devices [ECM98o]. Dialog [ECM00y, ECM00z].

Dictionary [ECM91a]. Digital

[ECM99q, ECM99r, ECM99s, ECM99t]. Disk [ECM88, ECM95a, ECM87b, ECM90d, ECM94a, ECM94b, ECM94c, ECM92j, ECM92k, ECM93g, ECM93h, ECM95b, ECM94d, ECM94e, ECM95f, ECM96h, ECM96i, ECM96j, ECM97-39, ECM99e, ECM99f, ECM99h, ECM99i, ECM98r, ECM98s, ECM00-32, ECM01d, ECM02a, ECM03b, ECM03c, ECM04f, ECM03i, ECM03j, ECM04g, ECM04m, ECM06e, ECM85b, ECM05c, ECM05i, ECM06f]. Disks [ECM96c, ECM98n, ECM04b]. Distribution [ECM96d]. Disturb [ECM97s, ECM97t]. Diversion [ECM97l, ECM97m, ECM00x, ECM04n]. DL [ECM05c, ECM06e]. DLT [ECM95h, ECM97-37, ECM97-38, ECM00m]. DLT1 [ECM92i]. DLT2 [ECM93j]. DLT3 [ECM94g]. DND [ECM97s, ECM97t]. Do [ECM97s, ECM97t, ECM97s]. DSS2 [ECM99q, ECM99r, ECM99s, ECM99t]. DTF [ECM98h, ECM00-30]. DTF-1 [ECM98h]. DTF-2 [ECM00-30]. Dual [ECM92c]. DVD [ECM02a, ECM03c, ECM04m, ECM99e, ECM99f, ECM99h, ECM98n, ECM98r, ECM04b]. DVD-R [ECM04m, ECM98r]. DVD-RAM [ECM02a, ECM99h, ECM98n, ECM04b]. DVD-Recordable [ECM98r]. DVD-RW [ECM03c].

E4X [ECM04k]. Each

[ECM88, ECM87b, ECM90d, ECM94e]. ECMA-100 [ECM88]. ECMA-106 [ECM93a]. ECMA-107 [ECM95a]. ECMA-108 [ECM96a]. ECMA-109 [ECM96b]. ECMA-113 [ECM99a]. ECMA-114 [ECM00a]. ECMA-118 [ECM86a]. ECMA-119 [ECM87a]. ECMA-120 [ECM93b]. ECMA-121 [ECM00b]. ECMA-125 [ECM87b]. ECMA-128 [ECM99b]. ECMA-13 [ECM85a]. ECMA-130 [ECM96c].

ECMA-133 [ECM98a]. ECMA-139 [ECM90a]. ECMA-142 [ECM97a]. ECMA-143 [ECM97b]. ECMA-144 [ECM00c]. ECMA-145 [ECM90b]. ECMA-146 [ECM90c]. ECMA-147 [ECM90d]. ECMA-148 [ECM97c]. ECMA-149 [ECM97d]. ECMA-150 [ECM92a]. ECMA-151 [ECM91a]. ECMA-152 [ECM93c]. ECMA-153 [ECM94a]. ECMA-154 [ECM94b]. ECMA-155 [ECM97e]. ECMA-156 [ECM93d]. ECMA-157 [ECM93e]. ECMA-158 [ECM97f]. ECMA-159 [ECM91b]. ECMA-160 [ECM92b]. ECMA-161 [ECM93f]. ECMA-162 [ECM97g]. ECMA-163 [ECM97h]. ECMA-164 [ECM97i]. ECMA-165 [ECM97j]. ECMA-167 [ECM97k]. ECMA-168 [ECM94c]. ECMA-169 [ECM92c]. ECMA-170 [ECM92d]. ECMA-171 [ECM92e]. ECMA-172 [ECM92f]. ECMA-173 [ECM97l]. ECMA-174 [ECM97m]. ECMA-175 [ECM98b]. ECMA-176 [ECM98c]. ECMA-177 [ECM97n]. ECMA-178 [ECM97o]. ECMA-179 [ECM92g]. ECMA-180 [ECM92h]. ECMA-182 [ECM92i]. ECMA-183 [ECM92j]. ECMA-184 [ECM92k]. ECMA-185 [ECM97p]. ECMA-186 [ECM00d]. ECMA-189 [ECM93g]. ECMA-190 [ECM93h]. ECMA-191 [ECM97q]. ECMA-192 [ECM97r]. ECMA-193 [ECM97s]. ECMA-194 [ECM97t]. ECMA-195 [ECM95b]. ECMA-196 [ECM93i]. ECMA-197 [ECM93j]. ECMA-198 [ECM95c]. ECMA-199 [ECM93k]. ECMA-200 [ECM93l]. ECMA-201 [ECM94d]. ECMA-202 [ECM97u]. ECMA-203 [ECM97v]. ECMA-205 [ECM93m]. ECMA-206 [ECM93n]. ECMA-207 [ECM94e]. ECMA-208 [ECM94f]. ECMA-209 [ECM94g]. ECMA-210 [ECM95d].

ECMA-211 [ECM97w]. ECMA-212 [ECM97x]. ECMA-213 [ECM97y]. ECMA-214 [ECM97z]. ECMA-215 [ECM97-27]. ECMA-216 [ECM97-28]. ECMA-217 [ECM94h]. ECMA-218 [ECM94i]. ECMA-219 [ECM96d]. ECMA-220 [ECM97-29]. ECMA-221 [ECM97-30]. ECMA-222 [ECM95e]. ECMA-223 [ECM95f]. ECMA-224 [ECM97-31]. ECMA-225 [ECM97-32]. ECMA-226 [ECM95g]. ECMA-230 [ECM97-33]. ECMA-231 [ECM95h]. ECMA-232 [ECM95i]. ECMA-233 [ECM97-34]. ECMA-234 [ECM95j]. ECMA-235 [ECM96e]. ECMA-236 [ECM96f]. ECMA-237 [ECM96g]. ECMA-238 [ECM96h]. ECMA-239 [ECM96i]. ECMA-240 [ECM96j]. ECMA-241 [ECM98d]. ECMA-242 [ECM98e]. ECMA-243 [ECM97-35]. ECMA-244 [ECM96k]. ECMA-245 [ECM97-36]. ECMA-246 [ECM98f]. ECMA-247 [ECM98g]. ECMA-248 [ECM98h]. ECMA-249 [ECM98i]. ECMA-250 [ECM98j]. ECMA-251 [ECM98k]. ECMA-252 [ECM96l]. ECMA-253 [ECM00e]. ECMA-254 [ECM99c]. ECMA-258 [ECM97-37]. ECMA-259 [ECM97-38]. ECMA-260 [ECM97-39]. ECMA-261 [ECM97-40]. ECMA-262 [ECM99d]. ECMA-263 [ECM98l]. ECMA-264 [ECM98m]. ECMA-265 [ECM97-41]. ECMA-266 [ECM97-42]. ECMA-267 [ECM99e]. ECMA-268 [ECM99f]. ECMA-269 [ECM00f]. ECMA-270 [ECM97-43]. ECMA-271 [ECM99g]. ECMA-272 [ECM99h]. ECMA-273 [ECM98n]. ECMA-274 [ECM99i]. ECMA-275 [ECM98o]. ECMA-276 [ECM98p]. ECMA-277 [ECM98q]. ECMA-278 [ECM00g]. ECMA-279 [ECM98r]. ECMA-280 [ECM98s]. ECMA-281 [ECM00h]. ECMA-282 [ECM00i].

ECMA-283 [ECM00j]. ECMA-284 [ECM00k]. ECMA-285 [ECM00l]. ECMA-286 [ECM00m]. ECMA-287 [ECM99j]. ECMA-288 [ECM99k]. ECMA-289 [ECM99l]. ECMA-290 [ECM99m]. ECMA-291 [ECM99n]. ECMA-292 [ECM99o]. ECMA-293 [ECM99p]. ECMA-294 [ECM99q]. ECMA-295 [ECM99r]. ECMA-296 [ECM99s]. ECMA-297 [ECM99t]. ECMA-298 [ECM99u]. ECMA-299 [ECM00n]. ECMA-300 [ECM00o]. ECMA-301 [ECM00p]. ECMA-302 [ECM00q]. ECMA-303 [ECM00r]. ECMA-304 [ECM00s]. ECMA-305 [ECM00t]. ECMA-306 [ECM00u]. ECMA-307 [ECM00v]. ECMA-308 [ECM00w]. ECMA-309 [ECM00x]. ECMA-310 [ECM00y]. ECMA-311 [ECM00z]. ECMA-312 [ECM00-27]. ECMA-313 [ECM00-28]. ECMA-314 [ECM00-29]. ECMA-315 [ECM00-30]. ECMA-316 [ECM00-31]. ECMA-317 [ECM00-32]. ECMA-318 [ECM00-33]. ECMA-319 [ECM01a]. ECMA-320 [ECM01b]. ECMA-321 [ECM01c]. ECMA-322 [ECM01d]. ECMA-323 [ECM04a]. ECMA-324 [ECM01e]. ECMA-325 [ECM01f]. ECMA-326 [ECM01g]. ECMA-327 [ECM01h]. ECMA-328 [ECM01i]. ECMA-329 [ECM01j]. ECMA-330 [ECM02a]. ECMA-331 [ECM04b]. ECMA-332 [ECM01k]. ECMA-333 [ECM03a]. ECMA-334 [ECM01l, ECM02b, ECM05a, ECM06a]. ECMA-335 [ECM01m, ECM02c, ECM05b, ECM06b]. ECMA-336 [ECM02d]. ECMA-337 [ECM03b]. ECMA-338 [ECM03c]. ECMA-339 [ECM02e]. ECMA-340 [ECM04c]. ECMA-341 [ECM04d]. ECMA-342 [ECM02f]. ECMA-343 [ECM03d]. ECMA-344 [ECM03e].

ECMA-345 [ECM03f]. ECMA-346 [ECM03g]. ECMA-347 [ECM03h]. ECMA-348 [ECM04e]. ECMA-349 [ECM04f]. ECMA-35 [ECM94j]. ECMA-350 [ECM03i]. ECMA-351 [ECM03j]. ECMA-352 [ECM03k]. ECMA-353 [ECM04g]. ECMA-354 [ECM04h]. ECMA-355 [ECM04i]. ECMA-356 [ECM04j]. ECMA-357 [ECM04k]. ECMA-358 [ECM04l]. ECMA-359 [ECM04m]. ECMA-360 [ECM04n]. ECMA-361 [ECM04o]. ECMA-362 [ECM04p]. ECMA-363 [ECM04q]. ECMA-364 [ECM05c]. ECMA-365 [ECM05d]. ECMA-366 [ECM05e]. ECMA-367 [ECM05f]. ECMA-368 [ECM05g, ECM05h]. ECMA-369 [ECM05h]. ECMA-370 [ECM06c]. ECMA-371 [ECM05i]. ECMA-372 [ECM05j]. ECMA-373 [ECM06d]. ECMA-374 [ECM06e]. ECMA-375 [ECM06f]. ECMA-376 [ECM06g]. ECMA-377 [ECM06h]. ECMA-43 [ECM91c]. ECMA-48 [ECM91d]. ECMA-6 [ECM91e]. ECMA-74 [ECM99v]. ECMA-94 [ECM86b]. ECMA-99 [ECM85b]. ECMAScript [ECM99d, ECM99m, ECM01h, ECM04k]. Eco [ECM06c]. Effect [ECM94a]. Effects [ECM96h, ECM98s, ECM00-32]. Eiffel [ECM05f]. Electric [ECM92f]. Electronic [ECM01i, ECM99j, ECM04d]. Embedded [ECM91a]. Emission [ECM96b]. Emissions [ECM92f, ECM01i, ECM04l]. Emitted [ECM96a, ECM99v]. Employment [ECM95g, ECM96k, ECM00e, ECM99l]. Emulation [ECM98q]. Environment [ECM99q, ECM99r, ECM97d, ECM97f, ECM97g, ECM97-33, ECM97-43]. Environmental [ECM04d]. Equipment [ECM96a, ECM96b, ECM92b, ECM93l, ECM96g, ECM01i, ECM99v, ECM99j, ECM95i]. Evaluation [ECM99g, ECM93m].

Exchange [ECM97b, ECM97j, ECM96l, ECM99c, ECM97-41, ECM97-42, ECM99q, ECM99r, ECM99s, ECM99t, ECM99u, ECM01f, ECM03e, ECM03h, ECM97i, ECM97m, ECM98c, ECM97o, ECM00d, ECM97r, ECM97t, ECM97v, ECM97x, ECM97z, ECM97-27, ECM97-28, ECM97-30, ECM97-32, ECM97-34, ECM98e, ECM97-35, ECM97-36, ECM98k, ECM98m, ECM00i, ECM00k, ECM00o, ECM00p, ECM00q, ECM00s, ECM00u, ECM00z, ECM00-29]. Exchanges [ECM98a]. Extended [ECM93c, ECM99g]. Extension [ECM98p, ECM94j].

Feature [ECM93f, ECM98b, ECM98c, ECM97-27, ECM97-29, ECM97-30, ECM97-31, ECM97-32, ECM97-34, ECM98j, ECM98k, ECM96l, ECM00p, ECM00q, ECM00-28, ECM00-29]. Features [ECM00j, ECM00k, ECM00r, ECM00s]. Field [ECM04c, ECM03k, ECM06d]. Fields [ECM92f, ECM93k]. File [ECM95a, ECM87a, ECM85a, ECM97k, ECM94c, ECM04q, ECM06g, ECM06h]. First [ECM03i]. FK [ECM93f]. Flexible [ECM88, ECM87b, ECM90d, ECM94e, ECM85b]. Flows [ECM97a, ECM97c, ECM97h, ECM97l, ECM98b, ECM97n, ECM97p, ECM97q, ECM97s, ECM97u, ECM97w, ECM97y, ECM97-29, ECM97-31, ECM98d, ECM98j, ECM98l, ECM00h, ECM00j, ECM00n, ECM00p, ECM00r, ECM00t, ECM00y, ECM00-28, ECM01e, ECM03d, ECM03g]. Format [ECM90a, ECM90c, ECM92a, ECM93c, ECM92c, ECM92d, ECM92e, ECM92i, ECM93j, ECM95c, ECM94f, ECM94g, ECM95d, ECM95h, ECM96f, ECM96i, ECM96j, ECM98f, ECM98g, ECM98h, ECM98i, ECM97-37, ECM97-38, ECM99i, ECM00g, ECM00m, ECM99k, ECM99n, ECM99o, ECM99p, ECM00-31, ECM01a, ECM01b, ECM03b, ECM04f, ECM05c, ECM05i, ECM06e, ECM97-43, ECM01j, ECM04q]. Formats [ECM06g, ECM06h]. Frequency [ECM96a, ECM93k]. ftprad [ECM88, ECM87b, ECM90d, ECM85b]. Functional [ECM97a, ECM97c, ECM97h, ECM97j, ECM97l, ECM98b, ECM97n, ECM97p, ECM97q, ECM97s, ECM97u, ECM97w, ECM97y, ECM97-29, ECM97-31, ECM98d, ECM98j, ECM99c, ECM98l, ECM00h, ECM00j, ECM00n, ECM00p, ECM00r, ECM00t, ECM00v, ECM00y, ECM00-28, ECM01e, ECM03d, ECM03g]. Functionality [ECM93m, ECM99g]. Functions [ECM96d, ECM95g, ECM96k, ECM00e, ECM99l, ECM03a, ECM02d, ECM91d].

Gbytes [ECM01d, ECM02a, ECM03b, ECM03c, ECM04f, ECM03i, ECM03j, ECM04g, ECM04m, ECM96h, ECM97-39, ECM99i, ECM98r, ECM98s, ECM00-32, ECM05c, ECM05d, ECM05i, ECM06e]. Generation [ECM03i]. Generic [ECM99c, ECM93d, ECM93f, ECM97j, ECM00v]. GF [ECM97j, ECM99c]. GHz [ECM04l]. Gigabyte [ECM92j]. Gigabytes [ECM92k, ECM95b]. Graphic [ECM99a, ECM00a, ECM86a, ECM00b, ECM99b, ECM86b]. Greek [ECM86a]. GSS [ECM96e]. GSS-API [ECM96e].

H.323 [ECM01g, ECM01k, ECM03a, ECM00v, ECM00w, ECM00x]. Handling [ECM00j, ECM00k, ECM00r, ECM00s]. Hebrew [ECM00b]. Helical [ECM90a, ECM90b, ECM90c, ECM92a, ECM92c, ECM92d, ECM92e, ECM95c, ECM95d, ECM96f, ECM98f, ECM98g, ECM98h, ECM98i, ECM99k, ECM99n, ECM99o, ECM99p, ECM00-30, ECM00-31, ECM01j]. HH [ECM98g]. HH-1 [ECM98g]. High [ECM96a, ECM05g]. High-Frequency [ECM96a]. HS [ECM96i, ECM05i]. HS-1 [ECM96i]. HVD [ECM06f]. HVD-ROM [ECM06f]. Hz [ECM92f].

ICT [ECM04l]. ID [ECM93e]. Identification [ECM00-28, ECM00-29, ECM03g, ECM03h, ECM97c, ECM93e, ECM97h, ECM97i]. IDL [ECM97-33]. II [ECM94h, ECM94i]. III [ECM00f, ECM00l, ECM04a, ECM04e]. Immunity [ECM93k, ECM93l, ECM96g]. Implementation [ECM99r, ECM99t]. including [ECM93n]. Incoming [ECM97-27]. Independent [ECM94f]. Indication [ECM98d, ECM98e]. induced [ECM98o]. Information [ECM95a, ECM96a, ECM96b, ECM87a, ECM85a, ECM90a, ECM97a, ECM90b, ECM90c, ECM97c, ECM92a, ECM91a, ECM94a, ECM91b, ECM97h, ECM97k, ECM94c, ECM92c, ECM92d, ECM92e, ECM97l, ECM98b, ECM97n, ECM97p, ECM93g, ECM93h, ECM97q, ECM97s, ECM95c, ECM93l, ECM97u, ECM95d, ECM97w, ECM97y, ECM97-29, ECM97-31, ECM96f, ECM96g, ECM98d, ECM98f, ECM98g, ECM98h, ECM98i, ECM98j, ECM98l, ECM00h, ECM00j, ECM99k, ECM99n, ECM99o, ECM99p, ECM00n, ECM00p, ECM00q, ECM00r, ECM00t, ECM00y, ECM00-28, ECM00-30, ECM00-31, ECM01e, ECM01j, ECM03d, ECM03g, ECM99v, ECM98j, ECM98k]. Infrastructure [ECM01m, ECM02c, ECM05b, ECM06b]. Integrated [ECM98a, ECM97a, ECM97b, ECM97c, ECM97e, ECM97h, ECM97i, ECM97j, ECM97l, ECM97m, ECM98b, ECM98c, ECM97n, ECM97o, ECM97p, ECM00d, ECM97q, ECM97r, ECM97s, ECM97t, ECM97u, ECM97v, ECM97w, ECM97x, ECM97y, ECM97z, ECM97-27, ECM97-28, ECM97-29, ECM97-30, ECM97-31, ECM97-32, ECM95g, ECM95i, ECM97-34, ECM98d, ECM98e, ECM97-35, ECM96k, ECM97-36, ECM98j, ECM98k, ECM96l, ECM00e, ECM99c, ECM97-40, ECM98l, ECM98m, ECM97-41, ECM97-42, ECM98p, ECM98q, ECM00h, ECM00i, ECM00j, ECM00k, ECM99l, ECM99u, ECM00n, ECM00o, ECM00p, ECM00q, ECM00r, ECM00s, ECM00t, ECM00u, ECM00y, ECM00z, ECM00-27, ECM00-28, ECM00-29, ECM00-33, ECM01e, ECM01f, ECM03a, ECM02d, ECM03d, ECM03e, ECM03f, ECM03g, ECM03h]. Intensity [ECM92b]. Inter [ECM97b, ECM97j, ECM95g, ECM96k, ECM96l, ECM99c, ECM97-41, ECM97-42, ECM99q, ECM99r, ECM99s, ECM99t, ECM99u, ECM01f, ECM03e, ECM03h, ECM97i, ECM97m, ECM98c, ECM97o, ECM00d, ECM97r, ECM97t, ECM97v, ECM97x, ECM97z, ECM97-27, ECM97-28, ECM97-30, ECM97-32, ECM97-34, ECM98e, ECM97-35, ECM97-36, ECM98k, ECM98m, ECM00i, ECM00k, ECM00o, ECM00q, ECM00s, ECM00u, ECM00z, ECM00-29]. Inter-Exchange [ECM97b, ECM97j, ECM96l, ECM99c, ECM97-41, ECM97-42, ECM99q, ECM99r, ECM99s, ECM99t, ECM99u, ECM01f, ECM03e, ECM03h, ECM97i, ECM97m, ECM98c, ECM97o, ECM00d, ECM97r, ECM97t, ECM97v, ECM97x, ECM97z, ECM97-27, ECM97-28, ECM97-30, ECM97-32, ECM97-34, ECM98e, ECM97-35, ECM97-36, ECM98k, ECM98m, ECM00i, ECM00k, ECM00o, ECM00q, ECM00s, ECM00u, ECM00z, ECM00-29]. Inter-PINX [ECM96k]. Inter-PTNX [ECM95g]. Interception [ECM97-29, ECM97-30]. Interchange [ECM88, ECM95a, ECM87a, ECM93b, ECM87b, ECM85a, ECM96c, ECM90a, ECM90b, ECM90c, ECM90d, ECM92a, ECM91a, ECM93c, ECM94a, ECM94b, ECM91b, ECM97k, ECM94c, ECM92c, ECM92d, ECM92e, ECM92i, ECM92j, ECM92k, ECM93g, ECM93h, ECM95b, ECM93i, ECM93j, ECM95c, ECM94d, ECM94e, ECM94g, ECM95d, ECM95f, ECM95h, ECM96f, ECM96h, ECM96i, ECM96j, ECM98f, ECM98g, ECM98h, ECM98i, ECM97-37, ECM97-38, ECM97-39, ECM97-43, ECM99i, ECM00g, ECM98s, ECM00m, ECM99k, ECM99n, ECM99o, ECM99p, ECM00-30, ECM00-31, ECM00-32, ECM01a, ECM01b, ECM01d, ECM01j, ECM03b, ECM04f, ECM03i, ECM03j, ECM04g, ECM05d, ECM05i, ECM06e, ECM85b, ECM05c]. Interconnect [ECM02f]. Interconnecting [ECM00-33]. Interface [ECM95j, ECM04c, ECM03k, ECM04j, ECM05h, ECM06d, ECM97-33]. Interruption [ECM98l, ECM98m, ECM98l, ECM98m]. Interworking [ECM00v, ECM00w, ECM00x, ECM01g, ECM01k, ECM02e, ECM04n, ECM04o]. Intrusion [ECM97u, ECM97v]. IP [ECM02d]. IP-QSIG [ECM02d]. Irreversible [ECM96h, ECM98s, ECM00-32]. ISDN [ECM99q, ECM99r, ECM99s, ECM99t]. ISO [ECM88, ECM87b, ECM90d, ECM94e]. ISSD [ECM97c].

kbit [ECM97a, ECM00e, ECM99l]. kbit/s [ECM97a, ECM00e, ECM99l]. Key [ECM93f, ECM96d]. Keypad [ECM93d]. kHz [ECM92f]. KP [ECM93d].

Labelling [ECM85a]. Language [ECM97f, ECM97g, ECM99d, ECM01l, ECM02b, ECM05a, ECM06a, ECM01m, ECM02c, ECM05b, ECM06b, ECM04e, ECM05j, ECM05f, ECM97-33]. Latin [ECM99a, ECM00a, ECM86a, ECM00b, ECM99b, ECM00c, ECM86b]. Latin/Arabic [ECM00a]. Latin/Cyrillic [ECM99a]. Latin/Greek [ECM86a]. Latin/Hebrew [ECM00b]. Layer [ECM97-41]. Length [ECM92a, ECM92d, ECM92e, ECM95c, ECM96f]. Levels [ECM92b]. Lightning [ECM93l]. Limits [ECM96g]. Lines [ECM98p]. Linkage [ECM00-28, ECM00-29]. Location [ECM97-28, ECM00p, ECM00q]. Lossless [ECM95e, ECM01c].

m [ECM92a, ECM92d, ECM92e, ECM95c, ECM96f]. M.O. [ECM94b]. MAC [ECM05g, ECM05h]. Magnetic [ECM93b, ECM85a, ECM90a, ECM90b, ECM90c, ECM92a, ECM93c, ECM92c, ECM92d, ECM92e, ECM92f, ECM92i, ECM93i, ECM93j, ECM95c, ECM93k, ECM94g, ECM95d, ECM95h, ECM96f, ECM98f, ECM98g, ECM98h, ECM98i, ECM97-37, ECM97-38, ECM00g, ECM00m, ECM99k, ECM99n, ECM99o, ECM99p, ECM00-30, ECM00-31, ECM01a, ECM01b, ECM01j]. Magneto [ECM94a, ECM01d]. Magneto-Optical [ECM94a, ECM01d]. Mailbox [ECM03g, ECM03h]. Make [ECM03d, ECM03e]. MammothTape [ECM99p]. MammothTape-2 [ECM99p]. Management [ECM93f, ECM93n]. Many [ECM03i, ECM96h, ECM98s, ECM00-32]. Mapping [ECM95g, ECM96k, ECM00e, ECM97-43, ECM99l, ECM03a, ECM02d, ECM95g]. MAPPING-CM-STATIC [ECM95g]. Mapping-UUS [ECM96k]. Mapping/16 [ECM00e]. Mapping/8 [ECM99l]. Mapping/IP [ECM02d]. Mapping/IP-QSIG [ECM02d]. Mbytes [ECM94e, ECM96j]. MCA [ECM03f]. MCM [ECM03g, ECM03h]. MCM-SD [ECM03g]. MCR [ECM03e]. MCRSD [ECM03d]. Measurement [ECM96a, ECM92f, ECM96g, ECM98o, ECM01i, ECM99v]. Measurements [ECM92b]. Mechanism [ECM96e]. Media [ECM97k, ECM94c]. Megabytes [ECM94d, ECM95f, ECM96i]. Message [ECM98e, ECM01e, ECM01f, ECM03f, ECM03g, ECM03h, ECM98d]. Method [ECM92b, ECM93g, ECM93h]. Methods [ECM96g, ECM04j, ECM04p]. MFM [ECM88, ECM87b, ECM90d, ECM85b]. MIC [ECM99n, ECM99o]. MID [ECM03g, ECM03h]. MID-SD [ECM03g]. mm [ECM88, ECM93b, ECM87b, ECM96c, ECM90a, ECM90b, ECM90c, ECM90d, ECM92a, ECM93c, ECM94a, ECM94b, ECM92c, ECM92d, ECM92e, ECM92i, ECM92j, ECM92k, ECM93g, ECM93h, ECM95b, ECM93i, ECM93j, ECM95c, ECM94d, ECM94e, ECM94g, ECM95d, ECM95f, ECM95h, ECM96f, ECM96h, ECM96i, ECM96j, ECM98f, ECM98g, ECM98h, ECM98i, ECM97-37, ECM97-38, ECM97-39, ECM99e, ECM99f, ECM99h, ECM98n, ECM99i, ECM00g, ECM98r, ECM98s, ECM00m, ECM99k, ECM99n, ECM99o, ECM99p, ECM00-30, ECM00-31, ECM00-32, ECM01a, ECM01b, ECM01d, ECM01j, ECM02a, ECM04b, ECM03b, ECM03c, ECM04f, ECM03i, ECM03j, ECM04g, ECM04m, ECM05c, ECM05d, ECM05i, ECM06e, ECM06f, ECM85b]. Mobility [ECM00h, ECM00i, ECM00j, ECM97-27, ECM97-28, ECM97-34, ECM97-35, ECM00k]. Mode [ECM93a, ECM97a, ECM97b, ECM95g, ECM96k, ECM00e, ECM99l]. Model [ECM97c, ECM97h, ECM97l, ECM98b, ECM97n, ECM97p, ECM97q, ECM97s, ECM97u, ECM97w, ECM97y, ECM97-29, ECM97-31, ECM98d, ECM98j, ECM98l, ECM00h, ECM00j, ECM00n, ECM00p, ECM00r, ECM00t, ECM00y, ECM00-28, ECM01e, ECM03d, ECM03g]. Monitoring [ECM03g, ECM03h]. moving [ECM98o]. Multiple [ECM94a, ECM93g, ECM93h]. Multiplexing [ECM99l, ECM00e]. MWI [ECM98e]. MWISD [ECM98d].

Name [ECM97i, ECM97h]. Near [ECM06d, ECM04c, ECM03k]. Network [ECM98b, ECM98c, ECM97-27, ECM97-29, ECM97-30, ECM97-31, ECM97-32, ECM97-34, ECM98j, ECM98k, ECM96l, ECM99c, ECM97-40, ECM97-41, ECM97-42, ECM00j, ECM00k, ECM99u, ECM00p, ECM00q, ECM00r, ECM00s, ECM00-28, ECM00-29, ECM00-33, ECM01e, ECM01f, ECM03a, ECM02d, ECM03d, ECM03e, ECM03f, ECM03g, ECM03h, ECM98a, ECM97a, ECM97b, ECM97c, ECM97h, ECM97i, ECM97j, ECM97l, ECM97m, ECM98b, ECM98c, ECM97n, ECM97o, ECM97p, ECM00d, ECM97q, ECM97r, ECM97s, ECM97t, ECM97u, ECM97v, ECM97w, ECM97x, ECM97y, ECM97z, ECM97-27, ECM97-28, ECM97-29, ECM97-30, ECM97-31, ECM97-32, ECM95g, ECM95i, ECM97-34, ECM98d, ECM98e, ECM97-35, ECM96k, ECM97-36, ECM98j, ECM98k, ECM00e, ECM98l, ECM98m, ECM98p, ECM98q, ECM00h, ECM00i, ECM00j, ECM00k, ECM99l, ECM00n, ECM00o, ECM00p, ECM00q, ECM00r, ECM00s, ECM00t, ECM00u, ECM00y, ECM00z, ECM00-27, ECM00-28, ECM00-29, ECM00-33]. Networks [ECM98q, ECM00v, ECM00w, ECM00x, ECM00-27, ECM01g, ECM01k, ECM02d, ECM02e, ECM04i, ECM04n, ECM04o, ECM93a, ECM97e, ECM93d, ECM93e, ECM93f, ECM03a]. NFC [ECM06d]. NFCIP [ECM04c, ECM03k, ECM04j, ECM04p]. NFCIP-1 [ECM04c, ECM04j, ECM04p]. NFCIP-2 [ECM03k]. NISD [ECM97h]. No [ECM99b, ECM00c, ECM99q, ECM99r, ECM99s, ECM99t, ECM86b]. Noise [ECM96a, ECM96b, ECM99v]. Non [ECM97k]. Non-Sequential [ECM97k].

O [ECM97s, ECM97t]. ODC [ECM05d]. Offer [ECM97q, ECM97r]. Office [ECM06h, ECM06g]. On-demand [ECM96k]. Once [ECM94a, ECM97k, ECM94c, ECM93g, ECM93h, ECM96h, ECM98s, ECM00-32, ECM03i]. Only [ECM94c, ECM99e, ECM99f, ECM05d, ECM96c, ECM94b]. Open [ECM06g, ECM06h]. Optical [ECM96c, ECM94a, ECM94b, ECM92j, ECM92k, ECM93g, ECM93h, ECM95b, ECM94d, ECM95f, ECM96h, ECM96i, ECM96j, ECM97-39, ECM99i, ECM98s, ECM00-32, ECM01d, ECM03b, ECM04f, ECM03i, ECM03j, ECM04g, ECM06e, ECM05c, ECM05i]. Oriented [ECM93m, ECM97-40, ECM99g]. Outgoing [ECM97-34]. Override [ECM97s, ECM97t].

P [ECM98l, ECM98m]. Paging [ECM95i]. Pair [ECM96k]. Parallel [ECM00g]. Part [ECM96d, ECM99q, ECM99r, ECM99s, ECM99t]. Path [ECM98c, ECM98b]. PCTE [ECM97d, ECM97f, ECM97g, ECM97-33, ECM97-43]. PD [ECM96j]. Phase [ECM96j, ECM97-39, ECM04a, ECM04e, ECM92g, ECM92h, ECM94h, ECM94i, ECM00f, ECM00l]. PHY [ECM05g, ECM05h]. PICS [ECM99r, ECM99t]. PINX [ECM98a, ECM96k, ECM97-36, ECM98p, ECM00-33]. PISN [ECM01e, ECM01f, ECM03a, ECM02d, ECM03d, ECM03e, ECM03g, ECM03h, ECM98a, ECM97a, ECM97b, ECM97c, ECM97h, ECM97i, ECM97j, ECM97l, ECM97m, ECM98b, ECM98c, ECM97n, ECM97o, ECM97p, ECM00d, ECM97q, ECM97r, ECM97s, ECM97t, ECM97u, ECM97v, ECM97w, ECM97x, ECM97y, ECM97z, ECM97-27, ECM97-28, ECM97-29, ECM97-30, ECM97-31, ECM97-32, ECM95g, ECM95i, ECM97-34, ECM98d, ECM98e, ECM97-35, ECM96k, ECM97-36, ECM98j, ECM98k, ECM96l, ECM00e, ECM99c, ECM97-40, ECM98l, ECM98m, ECM97-41, ECM97-42, ECM98p, ECM98q, ECM00h, ECM00i, ECM00j, ECM00k, ECM99l, ECM99q, ECM99r, ECM99s, ECM99t, ECM99u, ECM00n, ECM00o, ECM00p, ECM00q, ECM00r, ECM00s, ECM00t, ECM00u, ECM00y, ECM00z, ECM00-27, ECM00-28, ECM00-29, ECM00-33, ECM03f]. Point [ECM93a, ECM93d, ECM93e, ECM93f, ECM00-33]. Portable [ECM97d, ECM97f, ECM97g, ECM97-33, ECM97-43]. Power [ECM92b, ECM93k]. PR [ECM98c]. Prenegotiation [ECM99s, ECM99t]. Priority [ECM98l, ECM98m]. Private [ECM96l, ECM99c, ECM97-40, ECM97-41, ECM97-42, ECM00h, ECM00i, ECM00j, ECM00k, ECM99u, ECM03f, ECM93a, ECM98a, ECM97a, ECM97b, ECM97c, ECM97e, ECM93d, ECM93e, ECM93f, ECM97h, ECM97i, ECM97j, ECM97l, ECM97m, ECM98b, ECM98c, ECM97n, ECM97o, ECM97p, ECM00d, ECM97q, ECM97r, ECM97s, ECM97t, ECM97u, ECM97v, ECM97w, ECM97x, ECM97y, ECM97z, ECM97-27, ECM97-28, ECM97-29, ECM97-30, ECM97-31, ECM97-32, ECM95g, ECM95i, ECM97-34, ECM98d, ECM98e, ECM97-35, ECM96k, ECM97-36, ECM98j, ECM98k, ECM00e, ECM98l, ECM98m, ECM98p, ECM98q, ECM00h, ECM00i, ECM00j, ECM00k, ECM99l, ECM00n, ECM00o, ECM00p, ECM00q, ECM00r, ECM00s, ECM00t, ECM00u, ECM00y, ECM00z, ECM00-27, ECM00-28, ECM00-29, ECM00-33, ECM01e, ECM01f, ECM03a, ECM02d, ECM03d, ECM03e, ECM03g, ECM03h]. Privilege [ECM96d]. Procedure [ECM92f]. Procedures [ECM97b, ECM97j]. Product [ECM04l]. products [ECM04d]. Profile [ECM95i, ECM00-27, ECM01h, ECM03f]. Proforma [ECM99r, ECM99t]. Programming [ECM97f, ECM97g, ECM95j, ECM05f]. Protection [ECM98l, ECM98m]. Protocol [ECM93a, ECM97b, ECM93e, ECM97i, ECM97j, ECM97m, ECM98c, ECM92h, ECM00d, ECM97t, ECM97v, ECM97x, ECM97-27, ECM97-28, ECM94i, ECM97-30, ECM97-32, ECM97-34, ECM98e, ECM97-35, ECM97-36, ECM98k, ECM96l, ECM99c, ECM98m, ECM97-41, ECM97-42, ECM00k, ECM00l, ECM99q, ECM99s, ECM99u, ECM00q, ECM00s, ECM00u, ECM00-29, ECM04a, ECM01f, ECM04c, ECM03e, ECM03h, ECM03k, ECM04p, ECM93d, ECM93f, ECM97j, ECM97o, ECM97r, ECM97z, ECM00i, ECM99r, ECM99t, ECM00o, ECM00v, ECM00z]. Protocol-2 [ECM03k]. PRSD [ECM98b]. PSS1 [ECM00-27]. PTN [ECM93a, ECM93d, ECM93e, ECM93f]. PTNX [ECM95g]. PUM [ECM00h, ECM00i, ECM00j, ECM00k]. PUMCH [ECM00k]. PUMCHSD [ECM00j]. PUMR [ECM00i]. PUMRSD [ECM00h].

QSIG [ECM96l, ECM99c, ECM97-41, ECM97-42, ECM99u, ECM01f, ECM01g, ECM01k, ECM03a, ECM02d, ECM02e, ECM03e, ECM03h, ECM04i, ECM04n, ECM04o, ECM97b, ECM97i, ECM97j, ECM97m, ECM98c, ECM97o, ECM00d, ECM97r, ECM97t, ECM97v, ECM97x, ECM97z, ECM97-27, ECM97-28, ECM97-30, ECM97-32, ECM97-34, ECM98e, ECM97-35, ECM98k, ECM98m, ECM00i, ECM00k, ECM99q, ECM99r, ECM99s, ECM99t, ECM00o, ECM00q, ECM00s, ECM00u, ECM00v, ECM00w, ECM00x, ECM00z, ECM00-27, ECM00-29, ECM00-33, ECM03f]. QSIG-AOC [ECM97x]. QSIG-BC [ECM97b]. QSIG-CC [ECM00d]. QSIG-CF [ECM97m]. QSIG-CI [ECM97v]. QSIG-CIDL [ECM00-29]. QSIG-CINT [ECM97-30]. QSIG-CMN [ECM98k]. QSIG-CO [ECM97r]. QSIG-CPI [ECM98m]. QSIG-CT [ECM97o]. QSIG-CTAU [ECM97-35]. QSIG-CTLR [ECM97-28]. QSIG-CTMI [ECM97-27]. QSIG-CTMO [ECM97-34]. QSIG-DND [ECM97t]. QSIG-GF [ECM97j]. QSIG-MCM [ECM03h]. QSIG-MCR [ECM03e]. QSIG-MID [ECM03h]. QSIG-MWI [ECM98e]. QSIG-NA [ECM97i]. QSIG-PR [ECM98c]. QSIG-PUMCH [ECM00k]. QSIG-PUMR [ECM00i]. QSIG-RE [ECM97z]. QSIG-SD [ECM00z]. QSIG-SMS [ECM01f]. QSIG-SSCT [ECM00o]. QSIG-TC [ECM97-32]. QSIG-WTMAU [ECM00u]. QSIG-WTMCH [ECM00s]. QSIG-WTMLR [ECM00q].

R [ECM04f, ECM04m, ECM05c, ECM98r]. Radiated [ECM04l]. Radio [ECM95i]. RAM [ECM02a, ECM99h, ECM98n, ECM04b]. RapidIO [ECM02f]. Rate [ECM05g]. Re [ECM03c, ECM97z]. Re-recordable [ECM03c]. Read [ECM96c, ECM94a, ECM94b, ECM94c, ECM93g, ECM93h, ECM96h, ECM99e, ECM99f, ECM98s, ECM00-32, ECM03i, ECM05d]. Read-Only [ECM94c, ECM99e, ECM99f, ECM05d, ECM96c]. Recall [ECM97y, ECM97z]. Recordable [ECM04m, ECM03c, ECM98r]. Recording [ECM88, ECM87b, ECM90a, ECM90b, ECM90c, ECM90d, ECM92a, ECM97k, ECM92c, ECM92d, ECM92e, ECM95c, ECM95d, ECM96f, ECM98f, ECM98g, ECM98h, ECM98i, ECM99k, ECM99n, ECM99o, ECM99p, ECM00-30, ECM00-31, ECM01j, ECM03b, ECM04f, ECM05c, ECM06e, ECM85b, ECM05i]. Reference [ECM93a, ECM98a, ECM93d, ECM93e, ECM93f, ECM98p, ECM00-33]. Registration [ECM97-28, ECM00h, ECM00i, ECM00p, ECM00q]. Related [ECM96d]. Replacement [ECM98b, ECM98c]. Request [ECM03d, ECM03e]. RESD [ECM97y]. Rewritable [ECM94b, ECM97k, ECM99h, ECM02a, ECM03i]. RF [ECM04j]. ROM [ECM96c, ECM06f]. Rooms [ECM92b]. RPE [ECM95i]. Rules [ECM91c]. RW [ECM03b, ECM03c, ECM05i, ECM06e, ECM99i].

s [ECM97a, ECM00e, ECM99l, ECM93a, ECM93d, ECM93e, ECM93f]. SAAL [ECM97-41]. Safety [ECM99j]. SBC [ECM99u]. Scan [ECM90a, ECM90b, ECM90c, ECM92a, ECM92c, ECM92d, ECM92e, ECM95c, ECM95d, ECM96f, ECM98f, ECM98g, ECM98h, ECM98i, ECM99k, ECM99n, ECM99o, ECM99p, ECM00-30, ECM00-31, ECM01j]. Scanning [ECM92b]. SD [ECM03g, ECM97s, ECM98l, ECM00n, ECM00p, ECM00r, ECM00t, ECM00z]. SDLT1 [ECM01b]. SDSD [ECM00y]. Security [ECM93m, ECM93n, ECM96d, ECM99g]. Separated [ECM99q, ECM99r, ECM99u]. Sequential [ECM97k]. Serpentine [ECM00g]. Service [ECM97a, ECM97n, ECM97o, ECM97q, ECM97r, ECM97u, ECM97y, ECM97z, ECM97-28, ECM98d, ECM98e, ECM96k, ECM97-40, ECM00h, ECM00i, ECM00n, ECM00o, ECM00p, ECM00q, ECM00y, ECM00z, ECM01e, ECM01f, ECM03d, ECM03e, ECM97v]. Services [ECM93a, ECM98a, ECM97a, ECM97b, ECM97c, ECM97e, ECM93d, ECM93e, ECM93f, ECM97h, ECM97i, ECM97j, ECM97l, ECM97m, ECM98b, ECM98c, ECM97n, ECM97o, ECM92g, ECM97p, ECM00d, ECM97q, ECM97r, ECM97s, ECM97t, ECM97u, ECM97v, ECM97w, ECM97x, ECM97y, ECM97z, ECM97-27, ECM97-28, ECM94h, ECM97-29, ECM97-30, ECM97-31, ECM97-32, ECM95g, ECM95i, ECM97-34, ECM98d, ECM98e, ECM97-35, ECM96k, ECM97-36, ECM98j, ECM98k, ECM96l, ECM00e, ECM99c, ECM97-40, ECM98l, ECM98m, ECM97-41, ECM97-42, ECM00f, ECM98p, ECM98q, ECM00h, ECM00i, ECM00j, ECM00k, ECM99l, ECM99u, ECM00n, ECM00o, ECM00p, ECM00q, ECM00r, ECM00s, ECM00t, ECM00u, ECM00v, ECM00w, ECM00x, ECM00y, ECM00z, ECM00-27, ECM00-28, ECM00-29, ECM00-33, ECM01e, ECM01f, ECM03a, ECM02d, ECM03d, ECM03e, ECM03f, ECM03g, ECM03h, ECM04e, ECM05e, ECM01g, ECM01k, ECM02e, ECM04h, ECM05e]. Session [ECM04h, ECM05e, ECM05e]. Set [ECM91c, ECM91e]. Sets [ECM99a, ECM00a, ECM86a, ECM00b, ECM99b, ECM00c, ECM91d, ECM86b]. Short [ECM01e, ECM01f]. Side [ECM88, ECM87b, ECM90d, ECM94e, ECM04f, ECM05c, ECM06e, ECM98r, ECM02a, ECM03b, ECM03c, ECM04m, ECM05i]. Sides [ECM85b]. SIDF [ECM94f]. SIG [ECM97-36]. Signalling [ECM93a, ECM97b, ECM93d, ECM93e, ECM93f, ECM97i, ECM97j, ECM97m, ECM98c, ECM97o, ECM00d, ECM97r, ECM97t, ECM97v, ECM97x, ECM97z, ECM97-27, ECM97-28, ECM97-30, ECM97-32, ECM97-34, ECM98e, ECM97-35, ECM97-36, ECM98k, ECM96l, ECM99c, ECM98m, ECM97-41, ECM97-42, ECM00i, ECM00k, ECM99q, ECM99r, ECM99s, ECM99t, ECM99u, ECM00o, ECM00q, ECM00s, ECM00u, ECM00v, ECM00w, ECM00x, ECM00z, ECM00-29, ECM01f, ECM01g, ECM01k, ECM02e, ECM03e, ECM03h, ECM04n, ECM04o, ECM96k]. Simple [ECM00y, ECM00z]. Single [ECM99a, ECM00a, ECM86a, ECM00b, ECM99b, ECM00c, ECM86b, ECM00n, ECM00o]. Single-Byte [ECM99a, ECM00a, ECM86a, ECM00b, ECM99b, ECM00c, ECM86b]. SIP [ECM02e, ECM04n, ECM04o, ECM04i]. SLDC [ECM01c]. small [ECM98o]. SMS [ECM01f]. SMSSD [ECM01e]. Sound [ECM92b]. Specification [ECM97c, ECM97d, ECM97h, ECM97l, ECM98b, ECM97n, ECM97p, ECM97q, ECM97s, ECM97u, ECM97w, ECM97y, ECM97-29, ECM97-31, ECM98d, ECM98j, ECM99d, ECM98l, ECM98q, ECM00h, ECM00j, ECM99m, ECM99q, ECM99r, ECM99s, ECM99t, ECM00n, ECM00p, ECM00r, ECM00t, ECM00y, ECM00-28, ECM01e, ECM01l, ECM02b, ECM05a, ECM06a, ECM03d, ECM03g, ECM05j, ECM04k]. speed [ECM03b, ECM04f, ECM05c, ECM05i, ECM06e]. SS7 [ECM99q, ECM99r, ECM99s, ECM99t]. SSCT [ECM00n, ECM00o]. SSCT-SD [ECM00n]. SSF [ECM93g]. SSIG [ECM93a, ECM93d, ECM93e, ECM93f]. SSIG-BC [ECM93a]. SSIG-FK [ECM93f]. SSIG-ID [ECM93e]. SSIG-KP [ECM93d]. Standard [ECM95i, ECM00-27, ECM03f, ECM05g]. Statement [ECM99r, ECM99t]. STATIC [ECM95g]. Step [ECM00n, ECM00o]. Streaming [ECM01c]. Structure [ECM95a, ECM87a, ECM85a, ECM97k, ECM94c, ECM94j, ECM91c, ECM98o]. structure-borne [ECM98o]. Sub [ECM00e, ECM99l]. Sub-Multiplexing [ECM99l, ECM00e]. Subscriber [ECM99q, ECM99r, ECM99s, ECM99t]. Supplementary [ECM97c, ECM93d, ECM93e, ECM93f, ECM97h, ECM97i, ECM97j, ECM97l, ECM97m, ECM97n, ECM97o, ECM97p, ECM00d, ECM97q, ECM97r, ECM97s, ECM97t, ECM97u, ECM97v, ECM97w, ECM97x, ECM97y, ECM97z, ECM97-28, ECM98d, ECM98e, ECM97-35, ECM96k, ECM98l, ECM98m, ECM00h, ECM00i, ECM00n, ECM00o, ECM00p, ECM00q, ECM00t, ECM00u, ECM00v, ECM00w, ECM00x, ECM00y, ECM00z, ECM01g, ECM03d, ECM03e, ECM03g, ECM03h]. Support [ECM93d, ECM97j, ECM00v]. Supported [ECM92g, ECM92h, ECM94h, ECM94i, ECM00f, ECM00l, ECM04a]. Surges [ECM93l]. SYNC [ECM97-36]. SYNC-SIG [ECM97-36]. Synchronization [ECM97-36]. System [ECM99q, ECM99r, ECM99s, ECM99t, ECM94f]. System-Independent [ECM94f]. Systems [ECM02f].

Tape [ECM93b, ECM90a, ECM90b, ECM90c, ECM92a, ECM93c, ECM92c, ECM92d, ECM92e, ECM92i, ECM93i, ECM93j, ECM95c, ECM94g, ECM95d, ECM95h, ECM96f, ECM98f, ECM98g, ECM98h, ECM98i, ECM97-37, ECM97-38, ECM00g, ECM00m, ECM99k, ECM99n, ECM99o, ECM99p, ECM00-30, ECM00-31, ECM01a, ECM01b, ECM01j]. Tapes [ECM85a, ECM92a, ECM92d, ECM92e, ECM95c, ECM96f]. TC [ECM97-32, ECM96l]. TCSD [ECM97-31]. Techniques [ECM94j]. Technology [ECM96a, ECM96b, ECM93l, ECM96g, ECM97-39, ECM99v]. TED [ECM06c]. Telecommunication [ECM93a, ECM93d, ECM93e, ECM93f, ECM00v, ECM00w, ECM00x, ECM01g, ECM01k, ECM02e, ECM04i, ECM04n, ECM04o]. Telecommunications [ECM96a, ECM96b, ECM92g, ECM92h, ECM94h, ECM94i, ECM00f, ECM00l, ECM04a, ECM99v]. Terminal [ECM97-27, ECM97-28, ECM97-34, ECM97-35, ECM00p, ECM00q, ECM00r, ECM00s, ECM00t, ECM00u]. Test [ECM04j, ECM04p]. Tool [ECM97d, ECM97f, ECM97g, ECM97-33, ECM97-43]. Track [ECM93b, ECM93c, ECM92i, ECM93i, ECM93j, ECM94g, ECM95h, ECM97-37, ECM97-38, ECM00g, ECM00m, ECM01a, ECM01b]. Tracks [ECM88, ECM87b, ECM90d, ECM94e, ECM85b]. Traffic [ECM00-27]. Transfer [ECM97n, ECM97o, ECM00n, ECM00o, ECM00w, ECM04o]. Transit [ECM97-32, ECM97-31, ECM96l].

Tunnelling [ECM03a, ECM02d, ECM04i]. two [ECM99q, ECM99r, ECM99s, ECM99t]. Type [ECM88, ECM87b, ECM90d, ECM94a, ECM93g, ECM93h, ECM94e, ECM96h, ECM98s, ECM00-32].

UDO [ECM03i]. Ultra [ECM03i, ECM05g]. Ultrium [ECM01a]. Ultrium-1 [ECM01a]. UMDTM [ECM05d]. Universal [ECM04q]. Use [ECM00-27, ECM00-33, ECM03f]. User [ECM00h, ECM00i, ECM00j, ECM00k, ECM96k]. User-to-User [ECM96k]. Using [ECM88, ECM87b, ECM92b, ECM92d, ECM92e, ECM98s, ECM00-32, ECM85b, ECM90d, ECM92a, ECM94a, ECM97k, ECM93g, ECM93h, ECM95c, ECM96f, ECM96h, ECM96j, ECM97-39, ECM99i, ECM03b, ECM04f, ECM05c, ECM05i, ECM06e]. UUS [ECM96k].

Values [ECM96b]. VDUs [ECM92f, ECM93k]. vibration [ECM98o]. Volume [ECM95a, ECM87a, ECM97k, ECM94c]. VXA [ECM00-31]. VXA-1 [ECM00-31].

Waiting [ECM98d, ECM98e]. Web [ECM04e, ECM05e]. WI [ECM06d]. Wide [ECM90a, ECM90b, ECM90c, ECM92a, ECM92c, ECM92d, ECM92e, ECM95c, ECM95d, ECM96f, ECM98f, ECM98g, ECM98h, ECM98i, ECM99k, ECM99n, ECM99o, ECM99p, ECM00-30, ECM00-31, ECM01j]. Wideband [ECM05g]. Windows [ECM95j]. Wired [ECM06d]. Wireless [ECM00q, ECM00s, ECM00u, ECM00p, ECM00q, ECM00r, ECM00t]. WORM [ECM94a, ECM93g, ECM93h, ECM96h, ECM97-39, ECM98s, ECM00-32]. Write [ECM97k, ECM94c, ECM03i, ECM94a, ECM93g, ECM93h, ECM96h, ECM98s, ECM00-32]. Write-Once [ECM97k, ECM94c]. WS [ECM05e]. WS-Session [ECM05e]. WSDL [ECM04e].

WTMAU [ECM00t, ECM00u].

WTMAU-SD [ECM00t]. WTMCH [ECM00r, ECM00s]. WTMCH-SD [ECM00r]. WTMLR [ECM00p, ECM00q]. WTMLR-SD [ECM00p].

XML

[ECM04k, ECM04a, ECM06g, ECM06h]. XT [ECM97-37].

References

ECMA-13

[ECM85a] ECMA. ECMA-13: File Structure and Labelling of Magnetic Tapes for Information Interchange. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, fourth edition, December 1985. URL http:

                                                   //www.ecma.ch/ecma1/STAND/
                                                   ECMA-013.HTM.

ECMA-99

[ECM85b] ECMA. ECMA-99: Data Interchange on 130 mm Flexible Disk Cartridges Using MFM Recording at 13 262 ftprad on Both Sides, 3,8 Tracks per mm. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, September 1985. URL http:

                                                   //www.ecma.ch/ecma1/STAND/
                                                   ECMA-099.HTM.

ECMA-118

[ECM86a] ECMA. ECMA-118: 8-Bit Single-Byte Coded Graphic Character Sets — Latin/Greek Alphabet. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1986. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-118.HTM.

ECMA-94

[ECM86b] ECMA. ECMA-94: 8-Bit Single-Byte Coded Graphic Character Sets — Latin Alphabets No. 1 to No. 4. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1986. viii + 46 pp. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-094.HTM;   http:
//www.ecma-international.
org/publications/files/ECMA-
ST/Ecma-094.pdf.

ECMA-119

[ECM87a] ECMA. ECMA-119: Volume and File Structure of CDROM for Information Interchange. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1987. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-119.HTM.

ECMA-125

[ECM87b] ECMA. ECMA-125: Data Interchange on 90 mm Flexible Disk Cartridges Using MFM Recording at 15 916 ftprad on 80 Tracks on Each Side — ISO Type 302. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1987. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-125.HTM.

ECMA-100

[ECM88] ECMA. ECMA-100: Data Interchange on 90 mm Flexible Disk Cartridges Using MFM Recording at 7 958 ftprad on 80 Tracks on Each Side — ISO Type 301. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1988. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-100.HTM.

ECMA-139

[ECM90a] ECMA. ECMA-139: 3,81 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — DDS Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1990. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-139.HTM.

ECMA-145

[ECM90b] ECMA. ECMA-145: 8 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1990. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-145.HTM.

ECMA-146

[ECM90c] ECMA. ECMA-146: 3,81 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — DATA/DAT Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1990. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-146.HTM.

ECMA-147

[ECM90d] ECMA. ECMA-147: Data Interchange on 90 mm Flexible Disk Cartridges using MFM Recording at 31 831 ftprad on 80 Tracks on Each Side — ISO Type 303. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1990. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-147.HTM.

ECMA-151

[ECM91a] ECMA. ECMA-151: Data Compression for Information Interchange — Adaptive Coding with Embedded Dictionary — DCLZ Algorithm. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1991. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-151.HTM.

ECMA-159

[ECM91b] ECMA. ECMA-159: Data Compression for Information Interchange — Binary Arithmetic Coding Algorithm. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1991.

                                        URL http://www.ecma.ch/ecma1/
                                        STAND/ECMA-159.HTM.

ECMA-43

[ECM91c] ECMA. ECMA-43: 8-Bit Coded Character Set Structure and Rules. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, December 1991. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-043.HTM.

ECMA-48

[ECM91d] ECMA. ECMA-48: Control Functions for Coded Character Sets. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, fifth edition, June 1991. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-048.HTM.

ECMA-6

[ECM91e] ECMA. ECMA-6: 7-Bit Coded Character Set. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, sixth edition, December 1991. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-006.HTM.

ECMA-150

[ECM92a] ECMA. ECMA-150: 3,81 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — DDS-DC Format using 60 m and 90 m Length Tapes. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1992. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-150.HTM.

ECMA-160

[ECM92b] ECMA. ECMA-160: Determination of Sound Power Levels of Computer and Business Equipment Using Sound Intensity Measurements: Scanning Method in Controlled Rooms. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1992. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-160.HTM.

ECMA-169

[ECM92c] ECMA. ECMA-169: 8 mm Wide Magnetic Tape Cartridge Dual Azimuth Format for Information Interchange — Helical Scan Recording. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1992. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-169.HTM.

ECMA-170

[ECM92d] ECMA. ECMA-170: 3,81 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — DDS Format Using 60 m and 90 m Length Tapes. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1992. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-170.HTM.

ECMA-171

[ECM92e] ECMA. ECMA-171: 3,81 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — DATA/DAT-DC Format Using 60 m and 90 m Length Tapes. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1992. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-171.HTM.

ECMA-172

[ECM92f] ECMA. ECMA-172: Procedure for Measurement of Emissions of Electric and Magnetic Fields from VDUs from 5 Hz to 400 kHz. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1992. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-172.HTM.

ECMA-179

[ECM92g] ECMA. ECMA-179: Services for Computer Supported Telecommunications Applications (CSTA) Phase I. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1992. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-179.HTM.

ECMA-180

[ECM92h] ECMA. ECMA-180: Protocol for Computer Supported Telecommunications Applications (CSTA) Phase I. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1992. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-180.HTM.

ECMA-182

[ECM92i] ECMA. ECMA-182: Data Interchange on 12,7 mm 48-Track Magnetic Tape Cartridges — DLT1 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1992.

URL http://www.ecma.ch/ecma1/
STAND/ECMA-182.HTM.

ECMA-183

[ECM92j] ECMA. ECMA-183: Data Interchange on 130 mm Optical Disk Cartridges — Capacity: 1 Gigabyte per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1992. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-183.HTM.

ECMA-184

[ECM92k] ECMA. ECMA-184: Data Interchange on 130 mm Optical Disk Cartridges — Capacity: 1,3 Gigabytes per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1992. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-184.HTM.

ECMA-106

[ECM93a] ECMA. ECMA-106: Private Telecommunication Networks (PTN) — Signalling Protocol at the S Reference Point — Circuit Mode Basic Services (SSIG-BC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, December 1993. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-106.HTM.

ECMA-120

[ECM93b] ECMA. ECMA-120: Data Interchange on 12,7 mm 18-Track Magnetic Tape Cartridges. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, December 1993. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-120.HTM.

ECMA-152

[ECM93c] ECMA. ECMA-152: Data Interchange on 12,7 mm 18-Track Magnetic Tape Cartridges — Extended Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1993. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-152.HTM.

ECMA-156

[ECM93d] ECMA. ECMA-156: Private Telecommunication Networks (PTN) — Signalling at the S Reference Point — Generic Keypad Protocol for the Support of Supplementary Services (SSIG-KP). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1993. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-156.HTM.

ECMA-157

[ECM93e] ECMA. ECMA-157: Private Telecommunication Networks (PTN) — Signalling Protocol at the S Reference Point — Identification Supplementary Services (SSIG-ID). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1993. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-157.HTM.

ECMA-161

[ECM93f] ECMA. ECMA-161: Private Telecommunication Networks (PTN) — Signalling at the S Reference Point — Generic Feature Key Management Protocol for the Control of Supplementary Services (SSIG-FK). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1993. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-161.HTM.

ECMA-189

[ECM93g] ECMA. ECMA-189: Information Interchange on 300 mm Optical Disk Cartridges of the Write Once, Read Multiple (WORM) Type using the SSF Method. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1993. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-189.HTM.

ECMA-190

[ECM93h] ECMA. ECMA-190: Information Interchange on 300 mm Optical Disk Cartridges of the Write Once, Read Multiple (WORM) Type using the CCS Method. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1993. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-190.HTM.

ECMA-196

[ECM93i] ECMA. ECMA-196: Data Interchange on 12,7 mm 36-Track Magnetic Tape Cartridges. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1993. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-196.HTM.

ECMA-197

[ECM93j] ECMA. ECMA-197: Data Interchange on 12,7 mm 112-Track Magnetic Tape Cartridges — DLT2 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1993.

URL http://www.ecma.ch/ecma1/
STAND/ECMA-197.HTM.

ECMA-199

[ECM93k] ECMA. ECMA-199: Immunity of VDUs to Power Frequency Magnetic Fields. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1993.

URL http://www.ecma.ch/ecma1/
STAND/ECMA-199.HTM.

ECMA-200

[ECM93l] ECMA. ECMA-200: Immunity of Information Technology Equipment to Lightning Surges. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1993. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-200.HTM.

ECMA-205

[ECM93m] ECMA. ECMA-205: Commercially Oriented Functionality Class for Security Evaluation (COFC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1993. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-205.HTM.

ECMA-206

[ECM93n] ECMA. ECMA-206: Association Context Management including Security Context Management. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1993. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-206.HTM.

ECMA-153

[ECM94a] ECMA. ECMA-153: Information Interchange on 130 mm Optical Disk Cartridges of the Write Once, Read Multiple (WORM) Type, using the Magneto-Optical Effect. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1994. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-153.HTM.

ECMA-154

[ECM94b] ECMA. ECMA-154: Data Interchange on 90 mm Optical Disk Cartridges, Read only and Rewritable, M.O. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1994. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-154.HTM.

ECMA-168

[ECM94c] ECMA. ECMA-168: Volume and File Structure of Read-Only and Write-Once Compact Disk Media for Information Interchange. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1994. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-168.HTM.

ECMA-201

[ECM94d] ECMA. ECMA-201: Data Interchange on 90 mm Optical Disk Cartridges — Capacity: 230 Megabytes per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1994. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-201.HTM.

ECMA-207

[ECM94e] ECMA. ECMA-207: Data Interchange on 90 mm Flexible Disk Cartridges — 326 Data Tracks on each Side — Capacity: 21 Mbytes — ISO Type 305. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1994. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-207.HTM.

ECMA-208

[ECM94f] ECMA. ECMA-208: System-Independent Data Format — SIDF. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1994. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-208.HTM.

ECMA-209

[ECM94g] ECMA. ECMA-209: Data Interchange on 12,7 mm 128-Track Magnetic Tape Cartridges — DLT3 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1994.

URL http://www.ecma.ch/ecma1/
STAND/ECMA-209.HTM.

ECMA-217

[ECM94h] ECMA. ECMA-217: Services for Computer Supported Telecommunications Applications (CSTA) Phase II. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1994. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-217.HTM.

ECMA-218

[ECM94i] ECMA. ECMA-218: Protocol for Computer Supported Telecommunications Applications (CSTA) Phase II. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1994. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-218.HTM.

ECMA-35

[ECM94j] ECMA. ECMA-35: Character Code Structure and Extension Techniques. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, sixth edition, December 1994. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-035.HTM.

ECMA-107

[ECM95a] ECMA. ECMA-107: Volume and File Structure of Disk Cartridges for Information Interchange. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1995. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-107.HTM.

ECMA-195

[ECM95b] ECMA. ECMA-195: Data Interchange on 130 mm Optical Disk Cartridges — Capacity: 2 Gigabytes per Cartridge.

ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1995. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-195.HTM.

ECMA-198

[ECM95c] ECMA. ECMA-198: 3,81 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — DDS-2 Format using 120 m Length Tapes. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1995. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-198.HTM.

ECMA-210

[ECM95d] ECMA. ECMA-210: 12,65 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — DATA-D3-1 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1995. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-210.HTM.

ECMA-222

[ECM95e] ECMA. ECMA-222: Adaptive Lossless Data Compression Algorithm. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1995. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-222.HTM.

ECMA-223

[ECM95f] ECMA. ECMA-223: Data Interchange on 90 mm Optical Disk Cartridges — Capacity: 385 Megabytes per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1995. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-223.HTM.

ECMA-226

[ECM95g] ECMA. ECMA-226: Private Integrated Services Network (PISN) — Mapping Functions for the Employment of Dedicated Circuit Mode Connections as Inter-PTNX Connections (MAPPING-CM-STATIC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1995. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-226.HTM.

ECMA-231

[ECM95h] ECMA. ECMA-231: Data Interchange on 12,7 mm 128-Track Magnetic Tape Cartridges — DLT 4 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1995. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-231.HTM.

ECMA-232

[ECM95i] ECMA. ECMA-232: Private Integrated Services Network (PISN) — Profile Standard for the Connection of Radio Paging Equipment (RPE) to a PISN. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1995. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-232.HTM.

ECMA-234

[ECM95j] ECMA. ECMA-234: Application Programming Interface for Windows (APIW). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1995. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-234.HTM.

ECMA-108

[ECM96a] ECMA. ECMA-108: Measurement of High-Frequency Noise Emitted by Information Technology and Telecommunications Equipment. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, December 1996. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-108.HTM.

ECMA-109

[ECM96b] ECMA. ECMA-109: Declared Noise Emission Values of Information Technology and Telecommunications Equipment. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, fourth edition, December 1996. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-109.HTM.

ECMA-130

[ECM96c] ECMA. ECMA-130: Data Interchange on Read-only 120 mm Optical Data Disks (CD-ROM). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1996. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-130.HTM.

ECMA-219

[ECM96d] ECMA. ECMA-219: Authentication and Privilege Attribute Security Application with Related Key Distribution Functions — Part 1, 2 and 3. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, March 1996. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-219.HTM.

ECMA-235

[ECM96e] ECMA. ECMA-235: The ECMA GSS-API Mechanism. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, March 1996. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-235.HTM.

ECMA-236

[ECM96f] ECMA. ECMA-236: 3,81 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — DDS-3 Format using 125 m Length Tapes. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1996. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-236.HTM.

ECMA-237

[ECM96g] ECMA. ECMA-237: Limits and Methods of Measurement of Immunity Characteristics of Information Technology Equipment. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1996. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-237.HTM.

ECMA-238

[ECM96h] ECMA. ECMA-238: Data Interchange on 130 mm Optical Disk Cartridge of Type WORM (Write Once Read Many) using Irreversible Effects — Capacity: 2,6 Gbytes per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1996. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-238.HTM.

ECMA-239

[ECM96i] ECMA. ECMA-239: Data Interchange on 90 mm Optical Disk Cartridges — HS-1 Format — Capacity: 650 Megabytes per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1996. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-239.HTM.

ECMA-240

[ECM96j] ECMA. ECMA-240: Data Interchange on 120 mm Optical Disk Cartridges using Phase Change PD Format — Capacity: 650 Mbytes per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1996. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-240.HTM.

ECMA-244

[ECM96k] ECMA. ECMA-244: Private Integrated Services Network (PISN) — Mapping Functions for the Employment of a Circuit Mode Basic Service and the Supplementary Service User-to-User Signalling as a Pair of On-demand Inter-PINX Connections (Mapping-UUS). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1996. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-244.HTM.

ECMA-252

[ECM96l] ECMA. ECMA-252: Broadband Private Integrated Services Network (B-PISN) — Inter-Exchange Signalling Protocol — Transit Counter Additional Network Feature (B-QSIG-TC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1996. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-252.HTM.

ECMA-142

[ECM97a] ECMA. ECMA-142: Private Integrated Services Network (PISN) — Circuit Mode 64 kbit/s Bearer Services — Service Description, Functional Capabilities and Information Flows (BCSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-142.HTM.

ECMA-143

[ECM97b] ECMA. ECMA-143: Private Integrated Services Network (PISN) — Circuit Mode Bearer Services — Inter-Exchange Signalling Procedures and Protocol (QSIG-BC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June 1997. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-143.HTM.

ECMA-148

[ECM97c] ECMA. ECMA-148: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Identification Supplementary Services (ISSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-148.HTM.

ECMA-149

[ECM97d] ECMA. ECMA-149: Portable Common Tool Environment (PCTE) — Abstract Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, fourth edition, Decem-

ber 1997. URL http://www.
ecma.ch/ecma1/STAND/ECMA-
149.HTM.

ECMA-155

[ECM97e] ECMA. ECMA-155: Private Integrated Services Networks — Addressing. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-155.HTM.

ECMA-158

[ECM97f] ECMA. ECMA-158: Portable Common Tool Environment (PCTE) — C Programming Language Binding. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, fourth edition, December 1997. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-158.HTM.

ECMA-162

[ECM97g] ECMA. ECMA-162: Portable Common Tool Environment (PCTE) — Ada Programming Language Binding. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, fourth edition, December 1997. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-162.HTM.

ECMA-163

[ECM97h] ECMA. ECMA-163: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Name Identification Supplementary Services (NISD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, September 1997. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ECMA-163.HTM.

ECMA-164

[ECM97i] ECMA. ECMA-164: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Name Identification Supplementary Services (QSIG-NA). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, Septem-

ber 1997. URL http://www.
ecma.ch/ecma1/STAND/ECMA-
164.HTM.

ECMA-165

[ECM97j] ECMA. ECMA-165: Private Integrated Services Network (PISN) — Generic Functional Protocol for the Support of Supplementary Services — Inter-Exchange Signalling Procedures and Protocol (QSIG-GF). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-165.HTM.

ECMA-167

[ECM97k] ECMA. ECMA-167: Volume and File Structure for Write-Once and Rewritable Media using Non-Sequential Recording for Information Interchange. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-167.HTM.

ECMA-173

[ECM97l] ECMA. ECMA-173: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Call Diversion Supplementary Services (CFSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-173.HTM.

ECMA-174

[ECM97m] ECMA. ECMA-174: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Call Diversion Supplementary Services (QSIG-CF). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ECMA-174.HTM.

ECMA-177

[ECM97n] ECMA. ECMA-177: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Call Transfer Supplementary Service (CTSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, September 1997. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ECMA-177.HTM.

ECMA-178

[ECM97o] ECMA. ECMA-178: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Call Transfer Supplementary Service (QSIG-CT). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, September 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-178.HTM.

ECMA-185

[ECM97p] ECMA. ECMA-185: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Call Completion Supplementary Services (CCSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-185.HTM.

ECMA-191

[ECM97q] ECMA. ECMA-191: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Call Offer Supplementary Service (COSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-191.HTM.

ECMA-192

[ECM97r] ECMA. ECMA-192: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Call Offer Supplementary Service (QSIG-CO). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June 1997. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-192.HTM.

ECMA-193

[ECM97s] ECMA. ECMA-193: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Do Not Disturb and Do Not Disturb Override Supplementary Services (DND(O)SD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-193.HTM.

ECMA-194

[ECM97t] ECMA. ECMA-194: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Do Not Disturb and Do Not Disturb Override Supplementary Services (QSIG-DND(O)). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June

1997.  URL http://www.
ecma.ch/ecma1/STAND/ECMA-
194.HTM.

ECMA-202

[ECM97u] ECMA. ECMA-202: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Call Intrusion Supplementary Service (CISD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-202.HTM.

ECMA-203

[ECM97v] ECMA. ECMA-203: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Call Intrusion Supplementary Service (QSIG-CI). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June 1997. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-203.HTM.

ECMA-211

[ECM97w] ECMA. ECMA-211: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Advice of Charge Supplementary Services (AOCSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-211.HTM.

ECMA-212

[ECM97x] ECMA. ECMA-212: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Advice of Charge Supplementary Services (QSIG-AOC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-212.HTM.

ECMA-213

[ECM97y] ECMA. ECMA-213: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Recall Supplementary Service (RESD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-213.HTM.

ECMA-214

[ECM97z] ECMA. ECMA-214: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Recall Supplementary Service (QSIG-RE). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-214.HTM.

ECMA-215

[ECM97-27] ECMA. ECMA-215: Private Integrated Services Network (PISN) — Cordless Terminal Mobility (CTM) — Inter-Exchange Signalling Protocol — Cordless Terminal Incoming Call Additional Network Feature (QSIG-CTMI). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, September 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-215.HTM.

ECMA-216

[ECM97-28] ECMA. ECMA-216: Private Integrated Services Network (PISN) — Cordless Terminal Mobility (CTM) — Inter-Exchange Signalling Protocol — Cordless Terminal Location Registration Supplementary Service (QSIG-CTLR). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, September 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-216.HTM.

ECMA-220

[ECM97-29] ECMA. ECMA-220: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Call Interception Additional Network Feature (ANF-CINTSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ECMA-220.HTM.

ECMA-221

[ECM97-30] ECMA. ECMA-221: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Call Interception Additional Network Feature (QSIG-CINT). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-221.HTM.

ECMA-224

[ECM97-31] ECMA. ECMA-224: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Transit Counter Additional Network Feature (ANF-TCSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ECMA-224.HTM.

ECMA-225

[ECM97-32] ECMA. ECMA-225: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Transit Counter Additional Network Feature (QSIG-TC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-225.HTM.

ECMA-230

[ECM97-33] ECMA. ECMA-230: Portable Common Tool Environment (PCTE) — IDL Binding (Interface Definition Language). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-230.HTM.

ECMA-233

[ECM97-34] ECMA. ECMA-233: Private Integrated Services Network (PISN) — Cordless Terminal Mobility (CTM) — Inter-Exchange Signalling Protocol — Cordless Terminal Outgoing Call Additional Network Feature (QSIG-CTMO). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, September 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-233.HTM.

ECMA-243

[ECM97-35] ECMA. ECMA-243: Private Integrated Services Network (PISN) — Cordless Terminal Mobility (CTM) — Inter-Exchange Signalling Protocol — Cordless Terminal Authentication Supplementary Services (QSIG-CTAU). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, September 1997. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-243.HTM.

ECMA-245

[ECM97-36] ECMA. ECMA-245: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — PINX Clock Synchronization (SYNC-SIG). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, September 1997. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ECMA-245.HTM.

ECMA-258

[ECM97-37] ECMA. ECMA-258: Data Interchange on 12,7 mm 128-Track Magnetic Tape Cartridges — DLT 3-XT Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1997. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ECMA-258.HTM.

ECMA-259

[ECM97-38] ECMA. ECMA-259: Data Interchange on 12,7 mm 208-Track Magnetic Tape Cartridges — DLT 5 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1997. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-259.HTM.

ECMA-260

[ECM97-39] ECMA. ECMA-260: Data Interchange on 356 mm Optical Disk Cartridges — WORM, using Phase Change Technology Capacity: 14,8 and 25 Gbytes per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-260.HTM.

ECMA-261

[ECM97-40] ECMA. ECMA-261: Broadband Private Integrated Services Network (B-PISN) — Service Description — Broadband Connection Oriented Bearer Services (B-BCSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-261.HTM.

ECMA-265

[ECM97-41] ECMA. ECMA-265: Broadband Private Integrated Services Network (B-PISN) — Inter-Exchange Signalling Protocol — Signalling ATM Adaptation Layer (B-QSIG-SAAL).

ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, September 1997. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-265.HTM.

ECMA-266

[ECM97-42] ECMA. ECMA-266: Broadband Private Integrated Services Network (B-PISN) — Inter-Exchange Signalling Protocol — Basic Call/Connection Control (B-QSIG-BC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, September 1997.

URL http://www.ecma.ch/ecma1/
STAND/ECMA-266.HTM.

ECMA-270

[ECM97-43] ECMA. ECMA-270: Portable Common Tool Environment (PCTE) — Mapping from CASE Data Interchange Format (CDIF) to PCTE. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1997.

URL http://www.ecma.ch/ecma1/
STAND/ECMA-270.HTM.

ECMA-133

[ECM98a] ECMA. ECMA-133: Private Integrated Services Network (PISN) — Reference Configurations for PISN Exchanges (PINX). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1998. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-133.HTM.

ECMA-175

[ECM98b] ECMA. ECMA-175: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Path Replacement Additional Network Feature (ANF-PRSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, December 1998. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-175.HTM.

ECMA-176

[ECM98c] ECMA. ECMA-176: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Path Replacement Additional Network Feature (QSIG-PR). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, December 1998. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-176.HTM.

ECMA-241

[ECM98d] ECMA. ECMA-241: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Message Waiting Indication Supplementary Service (MWISD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, December 1998. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-241.HTM.

ECMA-242

[ECM98e] ECMA. ECMA-242: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Message Waiting Indication Supplementary Service (QSIG-MWI). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, December 1998. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-242.HTM.

ECMA-246

[ECM98f] ECMA. ECMA-246: 8 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — AIT-1 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1998. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ECMA-246.HTM.

ECMA-247

[ECM98g] ECMA. ECMA-247: 8 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — HH-1 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1998. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-247.HTM.

ECMA-248

[ECM98h] ECMA. ECMA-248: 12,65 mm Wide Magnetic Tape Cassette for Information Interchange — Helical Scan Recording — DTF-1 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1998. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-248.HTM.

ECMA-249

[ECM98i] ECMA. ECMA-249: 8 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — DA-2 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1998. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-249.HTM.

ECMA-250

[ECM98j] ECMA. ECMA-250: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Common Information Additional Network Feature (ANF-CMNSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1998. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-250.HTM.

ECMA-251

[ECM98k] ECMA. ECMA-251: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Common Information Additional Network Feature (QSIG-CMN). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1998. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-251.HTM.

ECMA-263

[ECM98l] ECMA. ECMA-263: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Call Priority Interruption and Call Priority Interruption Protection Supplementary Services (CPI(P)SD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1998. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-263.HTM.

ECMA-264

[ECM98m] ECMA. ECMA-264: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Call Priority Interruption and Call Priority Interruption Protection Supplementary Services (QSIG-CPI(P)). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1998. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-264.HTM.

ECMA-273

[ECM98n] ECMA. ECMA-273: Case for 120 mm DVD-RAM Disks. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, February 1998. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-273.HTM.

ECMA-275

[ECM98o] ECMA. ECMA-275: Measurement of structure-borne vibration induced by small air moving devices (AMDs). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1998. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-275.HTM.

ECMA-276

[ECM98p] ECMA. ECMA-276: Private Integrated Services Network (PISN) — Reference Configuration for PINX Extension Lines. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1998. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-276.HTM.

ECMA-277

[ECM98q] ECMA. ECMA-277: Private Integrated Services Network (PISN) — Circuit Emulation Specification — Emulation of Basic Access by ATM Networks. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1998. URL http://

                                        www.ecma.ch/ecma1/STAND/../
                                        STAND/ecma-277.HTM.

ECMA-279

[ECM98r] ECMA. ECMA-279: 80 mm (1,23 Gbytes per side) and 120 mm (3,95 Gbytes per side) DVD-Recordable Disk (DVD-R). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1998. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-279.HTM.

ECMA-280

[ECM98s] ECMA. ECMA-280: Data Interchange on 130 mm Optical Disk Cartridges of Type WORM (Write Once Read Many) Using Irreversible Effects — Capacity: 5,2 Gbytes per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1998. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-280.HTM.

ECMA-113

[ECM99a] ECMA. ECMA-113: 8-Bit Single-Byte Coded Graphic Character Sets — Latin/Cyrillic Alphabet. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-113.HTM.

ECMA-128

[ECM99b] ECMA. ECMA-128: 8-Bit Single-Byte Coded Graphic Character Sets — Latin Alphabet No.

5. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-128.HTM.

ECMA-254

[ECM99c] ECMA. ECMA-254: Broadband Private Integrated Services Network (B-PISN) — Inter-Exchange Signalling Protocol — Generic Functional Protocol (B-QSIG-GF). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-254.HTM.

ECMA-262

[ECM99d] ECMA. ECMA-262: EC-MAScript Language Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, Decem-

ber 1999. URL http://www.
ecma.ch/ecma1/STAND/ECMA-
262.HTM.

ECMA-267

[ECM99e] ECMA. ECMA-267: 120 mm DVD — Read-Only Disk. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-267.HTM.

ECMA-268

[ECM99f] ECMA. ECMA-268: 80 mm DVD — Read-Only Disk. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-268.HTM.

ECMA-271

[ECM99g] ECMA. ECMA-271: Extended Commercially Oriented Functionality Class for Security Evaluation (E — COFC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-271.HTM.

ECMA-272

[ECM99h] ECMA. ECMA-272: 120 mm DVD Rewritable Disk (DVD-RAM). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1999. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-272.HTM.

ECMA-274

[ECM99i] ECMA. ECMA-274: Data Interchange on 120 mm Optical Disk using +RW Format — Capacity: 3,0 Gbytes and 6,0 Gbytes. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-274.HTM.

ECMA-287

[ECM99j] ECMA. ECMA-287: Safety of electronic equipment. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1999. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-287.HTM.

ECMA-288

[ECM99k] ECMA. ECMA-288: 3,81 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — DDS-4 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-288.HTM.

ECMA-289

[ECM99l] ECMA. ECMA-289: Private Integrated Services Network (PISN) — Mapping Functions for the Employment of 64 kbit/s Circuit Mode Connections with 8 kbit/s Sub-Multiplexing (Mapping/8). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-289.HTM.

ECMA-290

[ECM99m] ECMA. ECMA-290: EC-MAScript Components Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-290.HTM.

ECMA-291

[ECM99n] ECMA. ECMA-291: 8 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording AIT-1 with MIC Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-291.HTM.

ECMA-292

[ECM99o] ECMA. ECMA-292: 8 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording AIT-2 with MIC Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-292.HTM.

ECMA-293

[ECM99p] ECMA. ECMA-293: 8 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — MammothTape-2 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-293.HTM.

ECMA-294

[ECM99q] ECMA. ECMA-294: B-ISDN and B-PISN — Digital Subscriber Signalling System No. two (DSS2), Broadband Inter-Exchange Signalling (B-QSIG), and Signalling System No. 7 (SS7) — Call Control in a Separated Call and Bearer Control Environment — Part 1: Protocol Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-294.HTM.

ECMA-295

[ECM99r] ECMA. ECMA-295: B-ISDN and B-PISN — Digital Subscriber Signalling System No. two (DSS2), Broadband Inter-Exchange Signalling (B-QSIG), and Signalling System No. 7 (SS7) — Call Control in a Separated Call and Bearer Control Environment — Part 2: Protocol Implementation Conformance Statement (PICS) Proforma Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-295.HTM.

ECMA-296

[ECM99s] ECMA. ECMA-296: B-ISDN and B-PISN — Digital Subscriber Signalling System No. two (DSS2), Broadband Inter-Exchange Signalling (B-QSIG), and Signalling System No. 7 (SS7) — Prenegotiation — Part 1: Protocol Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-296.HTM.

ECMA-297

[ECM99t] ECMA. ECMA-297: B-ISDN and B-PISN — Digital Subscriber Signalling System No. two (DSS2), Broadband Inter-Exchange Signalling (B-QSIG), and Signalling System No. 7 (SS7) — Prenegotiation — Part 2: Protocol Implementation Conformance Statement (PICS) Proforma Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-297.HTM.

ECMA-298

[ECM99u] ECMA. ECMA-298: Broadband Private Integrated Services Network (B-PISN) — Inter-Exchange Signalling Protocol — Separated Bearer Control (SBC) (B-QSIG-SBC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-298.HTM.

ECMA-74

[ECM99v] ECMA. ECMA-74: Measurement of Airborne Noise Emitted by Information Technology and Telecommunications Equipment. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, sixth edition, December 1999. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-074.HTM.

ECMA-114

[ECM00a] ECMA. ECMA-114: 8-Bit Single-Byte Coded Graphic Character Sets — Latin/Arabic Alphabet. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, Decem-

                                        ber 2000. URL http://www.
                                        ecma.ch/ecma1/STAND/ECMA-
                                        114.HTM.

ECMA-121

[ECM00b] ECMA. ECMA-121: 8-Bit Single-Byte Coded Graphic Character Sets — Latin/Hebrew Alphabet. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 2000. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-121.HTM.

ECMA-144

[ECM00c] ECMA. ECMA-144: 8-Bit Single-Byte Coded Character Sets — Latin Alphabet No. 6. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, December 2000. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-144.HTM.

ECMA-186

[ECM00d] ECMA. ECMA-186: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Call Completion Supplementary Services (QSIG-CC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, Febru-

ary 2000. URL http://www.
ecma.ch/ecma1/STAND/ECMA-
186.HTM.

ECMA-253

[ECM00e] ECMA. ECMA-253: Private Integrated Services Network (PISN) — Mapping Functions for the Employment of 64 kbit/s Circuit Mode Connection with 16 kbit/s Sub-multiplexing (Mapping/16). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, September 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-253.HTM.

ECMA-269

[ECM00f] ECMA. ECMA-269: Services for Computer Supported Telecommunications Applications (CSTA) Phase III. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, fourth edition, June 2000. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-269.HTM.

ECMA-278

[ECM00g] ECMA. ECMA-278: Data Interchange on 12,7 mm 128-Track Magnetic Tape Cartridge — Parallel Serpentine Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 2000. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-278.HTM.

ECMA-281

[ECM00h] ECMA. ECMA-281: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Private User Mobility (PUM) — Registration Supplementary Service (PUMRSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 2000. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-281.HTM.

ECMA-282

[ECM00i] ECMA. ECMA-282: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Private User Mobility (PUM) — Registration Supplementary Service (QSIG-PUMR). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 2000. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ECMA-282.HTM.

ECMA-283

[ECM00j] ECMA. ECMA-283: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Private User Mobility (PUM) — Call Handling Additional Network Features (PUM-CHSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 2000. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-283.HTM.

ECMA-284

[ECM00k] ECMA. ECMA-284: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Private User Mobility (PUM) — Call Handling Additional Network Features (QSIG-PUMCH). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-284.HTM.

ECMA-285

[ECM00l] ECMA. ECMA-285: Protocol for Computer Supported Telecommunications Applications (CSTA) Phase III. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 2000. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-285.HTM.

ECMA-286

[ECM00m] ECMA. ECMA-286: Data Interchange on 12,7 mm 208-Track Magnetic Tape Cartridges — DLT 6 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 2000. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-286.HTM.

ECMA-299

[ECM00n] ECMA. ECMA-299: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Single Step Call Transfer Supplementary Service (SSCT-SD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, February 2000. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-299.HTM.

ECMA-300

[ECM00o] ECMA. ECMA-300: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Single Step Call Transfer Supplementary Service (QSIG-SSCT). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, February 2000. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-300.HTM.

ECMA-301

[ECM00p] ECMA. ECMA-301: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Wireless Terminal Location Registration Supplementary Service and Wireless Terminal Information Exchange Additional Network Feature (WTMLR-SD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2000. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-301.HTM.

ECMA-302

[ECM00q] ECMA. ECMA-302: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Wireless Terminal Location Registration Supplementary Service and Wireless Terminal Information Exchange Additional Network Feature (QSIG-WTMLR). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-302.HTM.

ECMA-303

[ECM00r] ECMA. ECMA-303: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Wireless Terminal Call Handling Additional Network Features (WTMCH-SD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-303.HTM.

ECMA-304

[ECM00s] ECMA. ECMA-304: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Wireless Terminal Call Handling Additional Network Features (QSIG-WTMCH). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2000. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-304.HTM.

ECMA-305

[ECM00t] ECMA. ECMA-305: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Wireless Terminal Authentication Supplementary Services (WTMAU-SD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2000. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-305.HTM.

ECMA-306

[ECM00u] ECMA. ECMA-306: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Wireless Terminal Authentication Supplementary Services (QSIG-WTMAU). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-306.HTM.

ECMA-307

[ECM00v] ECMA. ECMA-307: Corporate Telecommunication Networks — Signalling Interworking between QSIG and H.323 — Generic Functional Protocol for the Support of Supplementary Services. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-307.HTM.

ECMA-308

[ECM00w] ECMA. ECMA-308: Corporate Telecommunication Networks — Signalling Interworking between QSIG and H.323 — Call Transfer Supplementary Services. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-308.HTM.

ECMA-309

[ECM00x] ECMA. ECMA-309: Corporate Telecommunication Networks — Signalling Interworking between QSIG and H.323 — Call Diversion Supplementary Services. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-309.HTM.

ECMA-310

[ECM00y] ECMA. ECMA-310: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Simple Dialog Supplementary Service (SDSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2000. URL

http://www.ecma.ch/ecma1/
STAND/ECMA-310.HTM.

ECMA-311

[ECM00z] ECMA. ECMA-311: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Simple Dialog Supplementary Service (QSIG-SD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-311.HTM.

ECMA-312

[ECM00-27] ECMA. ECMA-312: Private Integrated Services Network (PISN) — Profile Standard for the Use of PSS1 (QSIG) in Air Traffic Services Networks. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-312.HTM.

ECMA-313

[ECM00-28] ECMA. ECMA-313: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Call Identification and Call Linkage Additional Network Feature (CIDLSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, September 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-313.HTM.

ECMA-314

[ECM00-29] ECMA. ECMA-314: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Call Identification and Call Linkage Additional Network Feature (QSIG-CIDL). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, September 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-314.HTM.

ECMA-315

[ECM00-30] ECMA. ECMA-315: 12,65 mm Wide Magnetic Tape Cassette for Information Interchange — Helical Scan Recording — DTF-2. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2000. URL http:

//www.ecma.ch/ecma1/STAND/
ECMA-315.HTM.

ECMA-316

[ECM00-31] ECMA. ECMA-316: 8 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — VXA-1 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2000. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-316.HTM.

ECMA-317

[ECM00-32] ECMA. ECMA-317: Data Interchange on 300 mm Optical Disk Cartridges of Type WORM (Write Once Read Many) Using Irreversible Effects — Capacity: 30 Gbytes per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2000. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ECMA-317.HTM.

ECMA-318

[ECM00-33] ECMA. ECMA-318: Private Integrated Services Network (PISN) — Use of QSIG at the C Reference Point between a PINX and an Interconnecting Network. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2000. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ecma-318.htm.

ECMA-319

[ECM01a] ECMA. ECMA-319: Data Interchange on 12,7 mm — 384-Track Magnetic Tape Cartridges — Ultrium-1 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2001. URL

http://www.ecma.ch/ecma1/
STAND/ecma-319.htm.

ECMA-320

[ECM01b] ECMA. ECMA-320: Data Interchange on 12,7 mm — 448-Track Magnetic Tape Cartridges — SDLT1 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2001. URL

http://www.ecma.ch/ecma1/
STAND/ecma-320.htm.

ECMA-321

[ECM01c] ECMA. ECMA-321: Streaming Lossless Data Compression Algorithm — (SLDC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2001. URL

http://www.ecma.ch/ecma1/
STAND/ecma-321.htm.

ECMA-322

[ECM01d] ECMA. ECMA-322: Data Interchange on 130 mm Magneto-Optical Disk Cartridges — Capacity: 9,1 Gbytes per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2001. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-322.htm.

ECMA-324

[ECM01e] ECMA. ECMA-324: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Short Message Service (SMSSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2001. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-324.htm.

ECMA-325

[ECM01f] ECMA. ECMA-325: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Short Message Service (QSIG-SMS). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2001. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-325.htm.

ECMA-326

[ECM01g] ECMA. ECMA-326: Corporate Telecommunication Networks — Signalling Interworking between QSIG and H.323 — Call Completion Supplementary Services. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2001. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-326.htm.

ECMA-327

[ECM01h] ECMA. ECMA-327: EC-MAScript 3: Compact Profile. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June 2001. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-327.htm.

ECMA-328

[ECM01i] ECMA. ECMA-328: Detection and Measurement of Chemical Emissions from Electronic Equipment. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2001. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-328.htm.

ECMA-329

[ECM01j] ECMA. ECMA-329: 8 mm Wide Magnetic Tape Cartridge for Information Interchange — Helical Scan Recording — AIT-3 Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2001. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-329.htm.

ECMA-332

[ECM01k] ECMA. ECMA-332: Corporate Telecommunication Networks — Signalling Interworking between QSIG and H.323 — Basic Services. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2001. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-332.htm.

ECMA-334-1

[ECM01l] ECMA. ECMA-334: C# Language Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2001. ISBN ???? xiii + 479 pp. LCCN ???? URL http://

www.ecma-international.org/
publications/files/ecma-st/
ECMA-334.pdf;   http://www.
ecma-international.org/publications/
standards/Ecma-334.htm.

ECMA-335-1

[ECM01m] ECMA. ECMA-335: Common Language Infrastructure (CLI). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2001. xiii + 479 pp. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-335.htm.

ECMA-330

[ECM02a] ECMA. ECMA-330: 120 mm (4,7 Gbytes per side) and 80 mm (1,46 Gbytes per side) DVD Rewritable Disk (DVD-RAM). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 2002. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-330.htm. Also ISO/IEC

17592.

ECMA-334-2

[ECM02b] ECMA. ECMA-334: C# Language Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 2002. ISBN ???? xiv + 471 pp. LCCN ???? URL http:

//www.ecma-international.
org/publications/standards/
Ecma-334.htm;   http://www.
ecma-international.org/publications/
files/ecma-st/ECMA-334.pdf.

ECMA-335-2

[ECM02c] ECMA. ECMA-335: Common Language Infrastructure (CLI). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 2002. iv + 99 (Part I), vii + 164 (Part II), iv + 124 (Part III), i + 16 (Part IV), ii + 78 (Part V) pp. URL

http://www.ecma.ch/ecma1/
STAND/ecma-335.htm;   http:
//www.ecma-international.
org/publications/files/ECMA-
ST/Ecma-335.pdf;   http://
www.ecma-international.org/
publications/files/ECMA-ST/
Ecma-335-xml.zip.

ECMA-336

[ECM02d] ECMA. ECMA-336: Private Integrated Services Network (PISN) — Mapping Functions for the Tunnelling of QSIG through IP Networks (Mapping/IP-QSIG). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2002. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ecma-336.htm.

ECMA-339

[ECM02e] ECMA. ECMA-339: Corporate Telecommunication Networks — Signalling Interworking between QSIG and SIP — Basic Services. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2002. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ecma-339.htm.

ECMA-342

[ECM02f] ECMA. ECMA-342: RapidIO Interconnect Systems. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2002. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ecma-342.htm.

ECMA-333

[ECM03a] ECMA. ECMA-333: Private Integrated Services Network (PISN) — Mapping Functions for the Tunnelling of QSIG through H.323 Networks. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 2003. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-333.htm.

ECMA-337

[ECM03b] ECMA. ECMA-337: Data Interchange on 120 mm and 80 mm — Optical Disk using +RW Format — Capacity: 4,7 and 1,46 Gbytes per side (Recording speed up to 4X). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 2003. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-337.htm.

ECMA-338

[ECM03c] ECMA. ECMA-338: 80 mm (1,46 Gbytes per side) and 120 mm (4,70 Gbytes per side) DVD Re-recordable Disk (DVD-RW). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2003. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-338.htm.

ECMA-343

[ECM03d] ECMA. ECMA-343: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Make Call Request Supplementary Service (MCRSD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2003. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ecma-343.htm.

ECMA-344

[ECM03e] ECMA. ECMA-344: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Make Call Request Supplementary Service (QSIG-MCR). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2003. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ecma-344.htm.

ECMA-345

[ECM03f] ECMA. ECMA-345: Private Integrated Services Network (PISN) — Use of QSIG for Message Centre Access (MCA) Profile Standard. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2003. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ecma-345.htm.

ECMA-346

[ECM03g] ECMA. ECMA-346: Private Integrated Services Network (PISN) — Specification, Functional Model and Information Flows — Message Centre Monitoring and Mailbox Identification Supplementary Services (MCM-SD / MID-SD). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2003. URL

http://www.ecma.ch/ecma1/
STAND/ecma-346.htm.

ECMA-347

[ECM03h] ECMA. ECMA-347: Private Integrated Services Network (PISN) — Inter-Exchange Signalling Protocol — Message Centre Monitoring and Mailbox Identification Supplementary Services (QSIG-MCM / QSIG-MID). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2003. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-347.htm.

ECMA-350

[ECM03i] ECMA. ECMA-350: Data Interchange on 130 mm Rewritable and Write Once Read Many Ultra Density Optical (UDO) Disk Cartridges — Capacity: 30 Gbytes per Cartridge — First Generation. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2003. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-350.htm.

ECMA-351

[ECM03j] ECMA. ECMA-351: Data Interchange on 90 mm Optical Disk Cartridges — Capacity: 1,3 Gbytes per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2003. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ecma-351.htm.

ECMA-352

[ECM03k] ECMA. ECMA-352: Near Field Communication Interface and Protocol-2 (NFCIP-2). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2003. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ecma-352.htm.

ECMA-323

[ECM04a] ECMA. ECMA-323: XML Protocol for Computer Supported Telecommunications Applications (CSTA) Phase III. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June 2004. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ecma-323.htm.

ECMA-331

[ECM04b] ECMA. ECMA-331: Case for 120 mm and 80 mm DVD-RAM Disks. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, also iso/iec 17594. edition, June 2004. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ecma-331.htm.

ECMA-340

[ECM04c] ECMA. ECMA-340: Near Field Communication — Interface and Protocol (NFCIP-1). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2004. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-340.htm.

ECMA-341

[ECM04d] ECMA. ECMA-341: Environmental design considerations for electronic products. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 2004. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-341.htm.

ECMA-348

[ECM04e] ECMA. ECMA-348: Web Services Description Language (WSDL) for CSTA Phase III. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 2004. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-348.htm.

ECMA-349

[ECM04f] ECMA. ECMA-349: Data Interchange on 120 mm and 80 mm Optical Disk using +R Format — Capacity: 4,7 and 1,46 Gbytes per Side (Recording speed up to 8X). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, June 2004. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ecma-349.htm.

ECMA-353

[ECM04g] ECMA. ECMA-353: Data Interchange on 90 mm Optical Disk Cartridges — Capacity 2,3 Gbytes per Cartridge. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2004. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ecma-353.htm.

ECMA-354

[ECM04h] ECMA. ECMA-354: Application Session Services. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2004. URL

                                        http://www.ecma.ch/ecma1/
                                        STAND/ecma-354.htm.

ECMA-355

[ECM04i] ECMA. ECMA-355: Corporate Telecommunication Networks — Tunnelling of QSIG over SIP. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2004. URL http:

                                        //www.ecma.ch/ecma1/STAND/
                                        ecma-355.htm.

ECMA-356

[ECM04j] ECMA. ECMA-356: NFCIP-1 — RF Interface Test Methods. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2004. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-356.htm.

ECMA-357

[ECM04k] ECMA. ECMA-357: EC-MAScript for XML (E4X) Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2004. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-357.htm.

ECMA-358

[ECM04l] ECMA. ECMA-358: ICT Product Radiated Emissions: 1–6 GHz. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2004. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-358.htm.

ECMA-359

[ECM04m] ECMA. ECMA-359: 80 mm (1,46 Gbytes per side) and 120 mm (4,70 Gbytes per side) DVD Recordable Disk (DVD-R). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2004. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-359.htm.

ECMA-360

[ECM04n] ECMA. ECMA-360: Corporate telecommunication networks — Signalling interworking between QSIG and SIP — Call diversion. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2004. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-360.htm.

ECMA-361

[ECM04o] ECMA. ECMA-361: Corporate telecommunication networks — Signalling interworking between QSIG and SIP — Call transfer. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2004. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-361.htm.

ECMA-362

[ECM04p] ECMA. ECMA-362: NFCIP-1 — Protocol Test Methods. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2004. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-362.htm.

ECMA-363

[ECM04q] ECMA. ECMA-363: Universal 3D File Format. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2004.

URL http://www.ecma.ch/ecma1/
STAND/ecma-363.htm.

ECMA-334-3

[ECM05a] ECMA. ECMA-334: C# Language Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June 2005. ISBN ???? xvii + 523 pp. LCCN ???? URL http:

//www.ecma-international.
org/publications/standards/
Ecma-334.htm;   http://www.
ecma-international.org/publications/
files/ecma-st/ECMA-334.pdf.

ECMA-335-3

[ECM05b] ECMA. ECMA-335: Common Language Infrastructure (CLI). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June 2005. viii + 104 (Part I), viii + 192 (Part II), iv + 138 (Part III), ii + 20 (Part IV), i + 4 (Part V), ii + 60 (Part VI) pp. URL

http://www.ecma.ch/ecma1/
STAND/ecma-335.htm;   http:
//www.ecma-international.
org/publications/files/ECMA-
ST/Ecma-335.pdf;   http://
www.ecma-international.org/
publications/files/ECMA-ST/
Ecma-335-xml.zip.

ECMA-364

[ECM05c] ECMA. ECMA-364: Data interchange on 120 mm and 80 mm optical disk using +R DL Format — Capacity: 8,55 and 2,66 Gbytes per Side (Recording speed 2,4x). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2005. URL

http://www.ecma.ch/ecma1/
STAND/ecma-364.htm;   http:
//www.ecma-international.
org/publications/files/ECMA-
ST/ECMA-364.pdf.

ECMA-365

[ECM05d] ECMA. ECMA-365: Data Interchange on 60 mm Read-Only ODC — Capacity: 1,8 Gbytes (UMDTM). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2005. URL

http://www.ecma.ch/ecma1/
STAND/ecma-365.htm;   http:
//www.ecma-international.
org/publications/files/ECMA-
ST/ECMA-365.pdf.

ECMA-366

[ECM05e] ECMA. ECMA-366: WS-Session — Web Services for Application Session Services. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2005. URL

http://www.ecma.ch/ecma1/
STAND/ecma-366.htm;   http:
//www.ecma-international.
org/publications/files/ECMA-
ST/Ecma-366.pdf;   http://
www.ecma-international.org/
publications/files/ECMA-ST/
Ecma-366_ws-session.zip.

ECMA-367

[ECM05f] ECMA. ECMA-367: Eiffel analysis, design and programming language. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2005. URL http://

www.ecma-international.org/
publications/files/ECMA-ST/
ECMA-367.pdf;   http://www.
ecma-international.org/publications/
standards/Ecma-367.htm.

ECMA-368

[ECM05g] ECMA. ECMA-368: High Rate Ultra Wideband PHY and MAC Standard. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2005. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-368.htm.

ECMA-369

[ECM05h] ECMA. ECMA-369: MAC– PHY Interface for ECMA-

368. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2005. URL http://

www.ecma-international.org/
publications/files/ECMA-ST/
ECMA-369.pdf;   http://www.
ecma.ch/ecma1/STAND/ecma-
369.htm.

ECMA-371

[ECM05i] ECMA. ECMA-371: Data Interchange on 120 mm and 80 mm optical disk using +RW HS Format – Capacity: 4,7 and 1,46 Gbytes per side (recording speed 8x). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2005. URL http:

//www.ecma.ch/ecma1/STAND/
ecma-371.htm.

ECMA-372

[ECM05j] ECMA. ECMA-372: C++/ CLI Language Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2005. URL http://

www.ecma-international.org/
publications/files/ECMA-ST/
ECMA-372.pdf;   http://www.
ecma-international.org/publications/
standards/Ecma-372.htm.

ECMA-334-4

[ECM06a] ECMA. ECMA-334: C# Language Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, fourth edition, June 2006. ISBN ???? xix + 531 pp. LCCN ???? URL http:

//www.ecma-international.
org/publications/standards/
Ecma-334.htm;   http://www.
ecma-international.org/publications/
files/ecma-st/ECMA-334.pdf.

ECMA-335-4

[ECM06b] ECMA. ECMA-335: Common Language Infrastructure (CLI). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, fourth edition, June 2006. vii + 104 (Part I), viii + 191 (Part II), iv + 138 (Part III), ii + 20 (Part IV), i + 4 (Part V), ii + 57 (Part VI) pp. URL http:

//www.ecma-international.
org/publications/standards/
Ecma-335.htm;   http://www.
ecma-international.org/publications/
files/ECMA-ST/Ecma-335.pdf;
 http://www.ecma-international.
org/publications/files/ECMA-
ST/ECMA-335.zip.

ECMA-370

[ECM06c] ECMA. ECMA-370: TED — The Eco Declaration. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2006.

URL http://www.ecma-international.
org/publications/files/ECMA-
ST/ECMA-370.pdf;   http:/
/www.ecma.ch/ecma1/STAND/
ecma-370.htm.

ECMA-373

[ECM06d] ECMA. ECMA-373: Near Field Communication Wired Interface (NFC?WI). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, June 2006. URL http://

www.ecma-international.org/
publications/files/ECMA-ST/
ECMA-373.pdf;   http://www.
ecma-international.org/publications/
standards/Ecma-373.htm.

ECMA-374

[ECM06e] ECMA. ECMA-374: Data Interchange on 120 mm and 80 mm Optical Disk using +RW DL Format — Capacity: 8,55 and 2,66 Gbytes per Side (Recording speed 2,4x). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2006. URL http://

                                        www.ecma-international.org/
                                        publications/files/ECMA-ST/
                                        ECMA-374.pdf;   http://www.
                                        ecma-international.org/publications/
                                        standards/Ecma-374.htm.

ECMA-375

[ECM06f] ECMA. ECMA-375: Case for 120 mm HVD-ROM disk. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2006. URL http://

                                        www.ecma-international.org/
                                        publications/files/ECMA-ST/
                                        ECMA-375.pdf;   http://www.
                                        ecma-international.org/publications/
                                        standards/Ecma-375.htm.

ECMA-376

[ECM06g] ECMA. ECMA-376: Office Open XML File Formats. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2006. URL http:

                                        //www.ecma-international.
                                        org/news/TC45_current_work/
                                        OpenXML%20White%20Paper.pdf;
                                         http://www.ecma-international.

org/publications/files/ECMA-
ST/ECMA-376.pdf;   http://
www.ecma-international.org/
publications/files/ECMA-ST/
Office%20Open%20XML%20Part%
201%20(PDF).pdf;   http://
www.ecma-international.org/
publications/files/ECMA-ST/
Office%20Open%20XML%20Part%
202%20(PDF).pdf;   http://
www.ecma-international.org/
publications/files/ECMA-ST/
Office%20Open%20XML%20Part%
203%20(PDF).pdf;   http://
www.ecma-international.org/
publications/files/ECMA-ST/
Office%20Open%20XML%20Part%
204%20(PDF).pdf;   http://
www.ecma-international.org/
publications/files/ECMA-ST/
Office%20Open%20XML%20Part%
205%20(PDF).zip;   http://
www.ecma-international.org/
publications/standards/Ecma-
376.htm.

ECMA-377

[ECM06h] ECMA. ECMA-377: Office Open XML File Formats. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2006. URL http:

//www.ecma-international.
org/news/TC45_current_work/
OpenXML%20White%20Paper.pdf;
 http://www.ecma-international.
org/publications/files/ECMA-
ST/Office%20Open%20XML%20Part%
201%20(PDF).zip;   http://
www.ecma-international.org/
publications/files/ECMA-ST/
Office%20Open%20XML%20Part%

                                        202%20(PDF).zip;   http://
                                        www.ecma-international.org/
                                        publications/files/ECMA-ST/
                                        Office%20Open%20XML%20Part%
                                        203%20(PDF).zip;   http://
                                        www.ecma-international.org/
                                        publications/files/ECMA-ST/
                                        Office%20Open%20XML%20Part%
                                        204%20(PDF).zip;   http://
                                        www.ecma-international.org/
                                        publications/files/ECMA-ST/
                                        Office%20Open%20XML%20Part%
                                        205%20(PDF).zip;   http://
                                        www.ecma-international.org/
                                        publications/standards/Ecma-
                                        377.htm.